### v1

# Ravenea: A Benchmark for Multimodal Retrieval-Augmented Visual Culture Understanding

Abstract
As vision-language models (VLMs) become increasingly integrated into daily life, the need for accurate visual culture understanding is becoming critical. Yet, these models frequently fall short in interpreting cultural nuances effectively. Prior work has demonstrated the effectiveness of retrieval-augmented generation (RAG) in enhancing cultural understanding in text-only settings, while its application in multimodal scenarios remains underexplored. To bridge this gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a new benchmark designed to advance visual culture understanding through retrieval, focusing on two tasks: culture-focused visual question answering (cVQA) and culture-informed image captioning (cIC). Ravenea extends existing datasets by integrating over 10,000 Wikipedia documents curated and ranked by human annotators. With Ravenea, we train and evaluate seven multimodal retrievers for each image query, and measure the downstream impact of retrieval-augmented inputs across fourteen state-of-the-art VLMs. Our results show that lightweight VLMs, when augmented with culture-aware retrieval, outperform their non-augmented counterparts (by at least $3.2\%$ absolute on cVQA and $6.2\%$ absolute on cIC). This highlights the value of retrieval-augmented methods and culturally inclusive benchmarks for multimodal understanding.
1 Introduction
Vision-language models (VLMs) are increasingly deployed in real-world applications, from education to assistive technologies di2024educational ; de2024vqask ; bates-etal-2025-template ; yao2025countllm , where understanding not only visual content but also the surrounding cultural context is crucial. Despite achieving impressive performance on general tasks yue2024mmmu ; lin2014microsoft ; an2025multimodality ; an2025generalized , VLMs often struggle to capture cultural nuances, such as traditions, symbols, and region-specific practices that require external, culturally grounded knowledge (nayak-etal-2024-benchmarking, ; liu2025culturevlm, ; khanuja-etal-2024-image, ; cao2024exploring, ). For example, as shown in Figure 1, a VLM may incorrectly identify the season of a festival scene as ‘Autumn’, overlooking that the image depicts Kyoto’s Gion Festival, which occurs in July and corresponds to ‘Summer’. A promising approach to address this limitation is the integration of external knowledge through retrieval-augmented generation (RAG) lewis2020retrieval , which has shown success in improving cultural awareness in language models (seo2025valuesrag, ; lertvittayakumjorn2025towards, ). However, prior work in the culture domain has predominantly been confined to text-only settings. Meanwhile, existing culture-related multimodal datasets primarily focus on evaluating VLM outputs on culturally oriented tasks, with limited emphasis on the integration of external cultural knowledge. As a result, the potential of RAG to improve multimodal cultural understanding remains underexplored.
Figure 1: Effectiveness of culture-aware RAG. Given a culturally grounded visual question, VLMs enhanced with culture-aware RAG—retrieving relevant Wikipedia documents—generate more accurate answers than their non-RAG counterparts. (see Section 5.2).
To bridge the gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a manually curated dataset designed to evaluate cultural understanding in VLMs with retrieval support. We construct Ravenea based on two existing datasets: CVQA (romero2025cvqa, ), which includes culturally relevant visual questions and corresponding answers, and CCUB (liu2023towards, ), offering culturally contextualized captions to foster inclusivity in text-to-image generation  [^1] We reuse the cultural captions from CCUB as ground-truth references for the inverse task, image-to-text generation, specifically for culture-aware image captioning.. For each instance drawn from the source datasets, we append a set of Wikipedia documents that have been human-ranked based on their cultural relevance to the associated image. This curation effort, designed to ensure broad cultural representation, contains data related to eight countries and spans eleven diverse categories, comprising more than 1,800 images and 10,000 human-ranked documents. Ravenea thus provides a retrieval-augmented benchmark for evaluating cultural sensitivity in multimodal retrievers, and further allows for assessing how well VLMs integrate and make use of retrieved cultural context. Specifically, we focus on two culturally grounded tasks: (i) culture-focused visual question answering (cVQA) and (ii) culture-informed image captioning (cIC). We select these two tasks for their centrality in assessing cultural understanding in VLMs—question answering tests context-aware reasoning, while captioning evaluates generation sensitivity to cultural nuances.
With Ravenea, we first train and evaluate seven multimodal retrievers that use both visual and textual inputs to retrieve Wikipedia documents for a given query image based on their cultural relevance. Then, we evaluate a diverse set of state-of-the-art (SOTA) VLMs, including GPT-4.1 achiam2023gpt , LLaVA-OneVision-7B li2025llavaonevision , Pixtral agrawal2024pixtral , Phi-4 Multimodal abouelenin2025phi , the Gemma3 family team2025gemma , the Qwen2.5-VL family bai2025qwen2 , the InternVL3 family zhu2025internvl3 , and the Deepseek-VL2 family wu2024deepseek , each with and without multimodal retrieval, to assess the impact of retrieval augmentation on cultural understanding. Our dataset provides a testbed for assessing the cultural relevance capabilities of multimodal retrievers and the effectiveness of VLMs in consuming and using such retrieved cultural context.
Our contributions and key findings include:
  • Ravenea benchmark: We introduce Ravenea, the first benchmark aimed at evaluating VLMs and multimodal retrieval in leveraging external knowledge for visual culture understanding. It comprises a diverse, large-scale collection of human-curated, culturally related documents linked to images from eight countries across eleven categories, enabling evaluation on two tasks: culture-focused visual question answering and culture-informed image captioning. (Section 3)
  • Cultural grounding annotations enhance multi-modal retrieval: We evaluate seven retrievers that integrate visual and textual cues to retrieve culturally relevant documents. We find fine-tuning retrievers on culture-targeted annotations leads to marked gains in retrieval accuracy, highlighting the value of explicit cultural supervision. (Section 6.3)
  • Benefits of culture-aware retrieval: Culture-aware retrieval boosts task performance across VLMs, with lightweight models showing the greatest improvement. This suggests that such retrieval can seamlessly integrate into downstream VLM tasks, enhancing their performance. (Section 5.2)
  • Cross-cultural variation: Evaluation across eight countries reveals that VLMs exhibit distinct cultural preferences, with each model favoring different regional contexts—suggesting model-specific cultural biases. (Section 6.2)
Figure 2: Ravenea: A Multimodal Retrieval-Augmented Visual culturE uNdErstAnding dataset. Left: Examples of cVQA and cIC tasks. Middle: Geographic and categorical distribution of cultural references. Right: Performance comparison of 14 VLMs, evaluated with and without integration of our culture-aware retriever. Here, CaCLIP="culture-aware CLIP-L/14@224px ".
2 Related Work
Retrieval augmentation for cultural understanding. Retrieval augmentation has demonstrated significant efficacy for culture-related NLP tasks (conia-etal-2024-towards, ; hu-etal-2024-bridging, ; fung-etal-2023-normsage, ). Prior work uses sources like the World Values Survey for cultural question answering seo2025valuesrag ; yadav-etal-2025-beyond , or retrieves web and knowledge base content to enhance cultural contextualization lertvittayakumjorn2025towards . In multimodal settings, however, cultural retrieval remains underexplored. While some approaches retrieve culture-relevant images to fine-tune VLMs liu2025culturevlm , they require additional training. In contrast, our method introduces a plug-and-play retrieval system that enriches cultural grounding at inference time without modifying the base model.
Vision-language culture datasets. Several recent studies examine cultural understanding in VLMs through multicultural VQA liu-etal-2021-visually ; nayak-etal-2024-benchmarking , cuisine recognition winata-etal-2025-worldcuisines ; li-etal-2024-foodieqa , and concept-based image retrieval bhatia-etal-2024-local . Others benchmark cultural entity recognition using Wikipedia-based prompts nikandrou-etal-2025-crope or curate culture-specific image sets for value-based tasks yadav-etal-2025-beyond . While these efforts highlight current VLM limitations, they often rely on static, manually curated datasets. A natural extension to address these deficiencies is through culture-relevant multimodal retrieval—an area lacking dedicated datasets. Our work specifically aims to fill this critical research gap.
3 Ravenea Dataset
We construct our dataset by building culturally relevant document lists for each image in two existing culture-grounded datasets: CVQA romero2025cvqa , a widely used dataset for culture-focused VQA, and CCUB liu2023towards , a dataset designed to mitigate cultural bias in text-to-image generation with culturally contextualized captions. To ensure broad geographic and cultural coverage, we curate a diverse subset comprising images from seven countries in CVQA: China, Nigeria, Russia, Spain, Mexico, India, and Indonesia, and all five countries in CCUB: China, Korea, India, Mexico, and Nigeria (see Figure 2). After that, we separate the dataset construction process into three critical stages shown in Figure 3.
3.1 Dataset Construction
Data collection. The data collection process consists of two main steps: (i) culture-related captioning and (ii) document retrieval. For culture-related captioning, we generate culturally grounded captions for each image to facilitate more effective attachment of relevant documents. Since the CVQA lacks captions and the CCUB provides only brief descriptions, we employ GPT-4o to generate richer, culturally informative captions (see the prompt example in Table 10). For document retrieval, we first conduct a coarse filtering using the generated cultural captions as queries for a BM25 Robertson2009ThePR retriever to extract the semantically relevant documents from a large-scale corpus comprising over six million English Wikipedia documents  [^2] https://huggingface.co/datasets/wikimedia/wikipedia. To mitigate the impact of inaccurate captions and ensure precise document relevance, we then perform human annotation on the retrieved documents.
Figure 3: Ravenea construction pipeline. Left: A two-stage retrieval process to match each image with relevant documents. Middle: Decomposition of cultural relevance into three interpretable dimensions to improve human annotation. Right: Postprocessing methods for quality control.
Relevance annotation. Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each document provides meaningful background or contextual information that is relevant to the culture described in the caption or the image (see Appendix K). Specifically, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: Country association: Is the topic of the Wikipedia article associated with the same country as the image and its caption? Topic alignment: Does the topic of the Wikipedia article align with the semantic category of the image and its caption? Explicit visual representation: Is the topic of the Wikipedia article explicitly mentioned or visually represented in the image and its caption? Each dimension is framed as a binary (True / False) question to reduce ambiguity and improve annotation consistency. However, for the country association dimension (the first listed), we introduce an additional label, "Cannot be determined", to handle cases where this association is unclear from the annotator’s perspective. Additionally, annotators are also instructed to include the title and URL of any relevant Wikipedia article they believe is missing from the top-10 retrieved results. These manually suggested articles are treated as the most cultural references closely related to the given image (see details in Appendix G).
Quality control. To ensure the quality and consistency of our annotations, we implement several quality control methods. First of all, prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines.
Table 1: Statistics of the Ravenea dataset. The dataset is constructed by curating existing sources and augmenting them with over 10,000 wiki-derived documents to broaden cultural knowledge coverage and enhance content diversity.
Dataset Images Documents Pairs Questions Captions
CVQA 1,213 8,319 12,130 2,331 -
CCUB 655 4,441 6,550 - 655
Ravenea 1,868 11,580 18,680 2,331 655
To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks (see Figure 14). We also perform an additional quality check on a subset of the dataset. Specifically, for each selected countries, we employ an additional local quality checker who is tasked with manually reviewing the annotations to verify their accuracy and adherence to the guidelines. The quality checker reviews a random sample of annotated items, focusing on both the relevance labels and the justification behind any edge cases, such as borderline relevance or use of the “Cannot be determined” label. If inconsistencies or deviations from the annotation guidelines are identified, the affected samples are flagged for re-annotation. The overall acceptance rate from the meta quality checkers is 98.2%. The inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$artstein2017inter between the meta checker and annotator on the sampled annotations is 0.83.
Figure 4: Examples demonstrating the impact of CaCLIP Wikipedia retrieval integration on cVQA and cIC tasks using DeepseekVL2-Tiny. When augmented with culture-aware retrieval, the model exhibits enhanced sensitivity to cultural context.
3.2 Dataset Statistics
We present the statistics of our dataset in Table 1. The dataset comprises a total of 1,868 culturally diverse images, with approximately 65% originating from CVQA and the remaining 35% from CCUB. Each image is paired with a GPT-4o-generated cultural caption, as well as the top-10 ranked Wikipedia documents retrieved via a cultural relevance scoring pipeline (Figure 3), yielding a total of 18,680 image-document pairs (see illustrated examples in Figure 4). The collection spans eight countries and encompasses a broad spectrum of cultural domains, such as traditional attire, festivals, architecture, cuisine, and social practices (see more statistics in Appendix E).
4 Culture-aware Multimodal Retriever
Leveraging the Ravenea dataset, we train and evaluate seven multimodal retrievers to retrieve culturally relevant Wikipedia documents using both visual and textual inputs. We fine-tune five representative models—spanning both generative and discriminative paradigms—to optimize multimodal document retrieval. Performance is evaluated using standard retrieval metrics, including Mean Reciprocal Rank (MRR) shi2012climf , Precision@k (P@k) jarvelin2017ir , and Normalized Discounted Cumulative Gain (nDCG@k) wang2013theoretical , where $k\in\{1,3,5\}$. We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance. We fine-tune a VisualBERT-based li2019visualbert ; li-etal-2020-bert-vision reranker following standard BERT-style setups nogueira2019passage , and adapt two multimodal generators—VL-T5 cho2021unifying and LLaVA-OneVision-7B li2025llavaonevision —for end-to-end document retrieval chen2022corpusbrain ; yuan2024asking ; sun2023learning . To enhance cultural awareness in the contrastive retrieval, we introduce Culture-Aware Contrastive (CAC) learning, a supervised learning framework compatible with both CLIP and SigLIP architectures. We denote the culture-aware fine-tuned versions of CLIP-L/14@224px and SigLIP2-SO/14@384px using CAC as CaCLIP and CaSigLIP2, respectively.
4.1 Culture-aware Contrastive Learning
Given an image $I_{i}$ associated with $T$ textual descriptions $\{D_{i1},D_{i2},\dots,D_{iT}\}$, each document $D_{it}$ is annotated with a binary label $y_{it}\in\{0,1\}$, where $y_{it}=1$ indicates cultural relevance and $y_{it}=0$ indicates irrelevance. For each image–text pair $(I_{i},D_{it})$, we employ a shared vision-language encoder—such as CLIP—to obtain modality-specific representations: $\mathbf{E}_{I_{i}}=\mathcal{E}_{V}(I_{i})$ for the visual input and $\mathbf{E}_{D_{it}}=\mathcal{E}_{L}(D_{it})$ for the textual input. We then compute the cosine similarity score $s_{it}$ between $\mathbf{E}_{I_{i}}$ and each corresponding $\mathbf{E}_{D_{it}}$, resulting in a similarity vector $\mathbf{S}_{i}=[s_{i1},s_{i2},\dots,s_{iT}]$.
Culture-awareness classification now amounts to:
$\mathcal{L}_{\text{Culture Classify}}=-\frac{1}{B\cdot T}\sum_{i=1}^{B}\sum_{j=1}^{T}\left[y_{ij}\log\sigma(s_{ij})+(1-y_{ij})\log(1-\sigma(s_{ij}))\right],$ (1)
where B is the number of images; $\sigma(\cdot)$ denotes the sigmoid function.
To prioritize culturally relevant descriptions in the ranking, we apply a margin ranking loss between all pairs of descriptions with differing cultural relevance. For each image $I_{i}$, we compare all pairs $(D_{ij},D_{ik})$ such that $y_{ij}=1$ and $y_{ik}=0$, and encourage the model to assign a higher similarity score to the relevant description. The ranking loss is defined as:
$\mathcal{L}_{\text{Rank}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{\begin{subarray}{c}j,k=1\\ y_{ij}=1,y_{ik}=0\end{subarray}}^{T}\max\left(0,\delta-(s_{ij}-s_{ik})\right),$ (2)
To mitigate the risk of overly similar positive text embeddings for the same image, we introduce a penalty that encourages intra-modal diversity among textual representations. We apply a diversity-promoting loss that forces the similarity between different text embeddings to be reduced while keeping each embedding highly similar to itself. Specifically, the penalty is formulated using an exponential function to emphasize the dissimilarity between embeddings:
$\mathcal{L}_{\text{Diversity}}\left(\mathbf{S}_{i}\right)=-\sum_{t=1}^{T}\log\left(\frac{\exp(s_{it})}{\sum_{j=1}^{T}\exp(s_{ij})}\right)$ (3)
Then we can get the culture-aware contrastive loss:
$\mathcal{L}_{\text{CAC}}=\frac{1}{3}\left(\mathcal{L}_{\text{Culture Classify}}+\mathcal{L}_{\text{Rank}}+\mathcal{L}_{\text{Diversity}}\right).$ (4)
Table 2: Performance with different retriever models. Fine-tuned contrastive models consistently outperform their frozen counterparts across tasks. Here, "CaSigLIP"=Culture-aware SigLIP2-SO/14@384px, and "CaCLIP"=Culture-aware CLIP-L/14@224px. Models in gray are frozen.
Method MRR $\uparrow$ P@1 $\uparrow$ P@3 $\uparrow$ P@5 $\uparrow$ nDCG@1 $\uparrow$ nDCG@3 $\uparrow$ nDCG@5 $\uparrow$ SigLIP2-SO/14@384px 66.71 50.42 40.14 34.25 58.85 62.43 67.50 CLIP-L/14@224px 70.76 54.58 43.47 36.58 62.54 67.25 72.43 VisualBERT 59.66 42.50 35.97 32.75 51.29 55.49 62.29 VL-T5 55.53 35.42 34.03 31.42 45.21 53.18 59.62 LLaVA-OneVision-7B 54.15 36.20 31.83 28.69 45.85 50.82 56.74 CaSigLIP2 (ours) 69.35 54.17 43.47 36.50 61.98 66.75 71.98 CaCLIP (ours) 78.34 65.42 49.44 39.50 72.25 75.22 79.32
4.2 Multimodal Retrieval Results
We perform a comprehensive evaluation of both frozen and fine-tuned retrievers, and present the results in Table 2. We find that fine-tuned models, particularly those based on contrastive learning, consistently outperform their frozen counterparts. For instance, CaCLIP achieves a substantial improvement in P@1, rising from 54.58% to 65.42%, and sets a new SOTA across all evaluation metrics. Although SigLIP2-SO/14@384px also benefits from fine-tuning, the performance gains are comparatively modest. In contrast, models such as LLaVA-OneVision-7B, VL-T5, and VisualBERT lag behind after fine-tuning, even underperforming relative to frozen baselines. This underperformance likely stems from the fact that models such as LLaVA-OneVision-7B and VisualBERT were originally pretrained for generative tasks with different objectives, whereas CLIP-L/14@224px and SigLIP2-SO/14@384px were explicitly trained for similarity-based alignment, providing them with a structural advantage in retrieval settings.
5 Multimodal Retrieval-augmented Visual Culture Understanding
We then evaluate the effectiveness of these retrievers with 14 SOTA VLMs, spanning a diverse set of architectures. We conduct experiments on two downstream tasks: cVQA and cIC, respectively.
5.1 Experimental Setup
Models.
We benchmark open and closed-weight SOTA VLMs on Ravenea, leveraging various retrievers against non-RAG baselines, assessing retrieval effectiveness across models of different sizes. The open-weight models include LLaVA-OneVision-7B li2025llavaonevision , Pixtral-12B agrawal2024pixtral , Phi-4 Multimodal-Instruct abouelenin2025phi , Gemma3-4B-Instruct and 27B-Instruct team2025gemma , Qwen2.5-VL-Instruct (3B, 7B, 72B  [^3] Due to computational constraints, we use the quantized version Qwen2.5-VL-Instruct-72B-AWQ.bai2025qwen2 , InternVL3 (2B, 8B, 38B) zhu2025internvl3 , and Deepseek-VL2 variants (Tiny and Base) wu2024deepseek . For the closed models, we adopt GPT-4.1 achiam2023gpt (accessed on 2025/04/14)  [^4] Knowledge cutoff: June 1, 2024; https://platform.openai.com/docs/models/gpt-4.1.
Evaluation metrics.
Table 3: Kendall’s $\tau$ rank correlation kendall1938new between automatic metrics and human judgments for the CCUB task. Statistically significant correlations ($p<0.05$) are marked with ✓. Our proposed metrics correlate stronger with human evaluation than the others.
Rouge-L lin-2004-rouge CIDER vedantam2015cider BERTScore Zhang*2020BERTScore: CLIPScore hessel-etal-2021-clipscore RegionScore (ours)
-0.172 ✗ -0.316 ✗ -0.011 ✗ 0.139 ✗ 0.442 ✓
For the cVQA task, we use accuracy as the primary evaluation metric, which measures the proportion of correctly predicted answers. For the cIC task, we employ several evaluation metrics including ROUGE-L lin-2004-rouge , CIDEr vedantam2015cider , BERTScore Zhang*2020BERTScore: , and CLIPScore hessel-etal-2021-clipscore , to assess the alignment between generated and reference captions across lexical, syntactic, and embedding-based levels. To further evaluate the cultural relevance and human-perceived quality, we conduct a human evaluation study. We employed four researchers to select the most accurate caption from 14 VLMs (see details in Appendix I and L), and find a significant mismatch between automatic metric scores and human judgments of cultural appropriateness (see Table 3). To bridge this gap, we further introduce RegionScore, a novel evaluation metric designed to quantify cultural grounding (see details in Appendix H and Table 9). It measures how well captions identify the correct country names tied to cultural elements, adding geographic and cultural specificity in image captioning.
5.2 Overall Performance
We present the main results in Table 4. The results demonstrate the efficacy of incorporating culture-aware retrieval augmentation. Employing fine-tuned retrievers yields substantial performance gains over both non-RAG and frozen retrievers baselines. Specifically, the CaCLIP achieves the highest average performance across both tasks, improves the accuracy for cVQA from 67.7% to 71.5%, and substantially improving the RegionScore for cIC from 40.2% to 58.1%. While CLIP-L/14@224px also offers improvements, fine-tuning consistently unlocks further potential. Furthermore, in the cVQA task, among all evaluated models, GPT-4.1 achieves the highest accuracy (86.8%) without RAG. Within the category of open-weight models, Qwen2.5-VL-72B leads with an accuracy of 81.0%. For lightweight models ($\leq$ 8B parameters), Qwen2.5-VL-7B achieves the best performance without RAG, reaching 67.7%. However, incorporating a CaCLIP significantly boosts performance, which enables InternVL3-8B to achieve 74.2% and outperform Qwen2.5-VL-7B by 0.6% with identical reranking. Notably, across multiple model families, augmenting the smallest variant with CaCLIP consistently elevates its performance to match or even exceed that of the next larger model tier. In the cIC task, with culture-aware contrastive learning, CaCLIP demonstrates substantial gains in identifying the culture in country-level of visual content, especially when built on top of VLMs with strong vision-language priors. It achieves the highest average RegionScore (58.1%) among the six reranking methods evaluated, with peak performance reaching 76.3% on the Gemma3-4B backbone. CaCLIP achieves leading the scores on 9 of 14 diverse suits of VLMs. This result underscores CaCLIP’s robustness and adaptability, particularly in culture-aware image captioning and retrieval tasks that demand fine-grained multimodal alignment.
Table 4: cVQA and cIC Performance w/ and w/o RAG. Models in gray are frozen retrievers. Results are colored as Best. VLMs augmented with finetuned retriever generally perform better.
Open Weights Closed Weights Retriever Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $cVQA_{Accuracy\uparrow}$ W/O RAG 67.7 57.1 65.5 64.2 67.7 81.0 58.4 66.5 71.9 63.6 78.7 55.5 66.1 64.8 86.8 SigLIP2-SO/14@384px 67.2 58.7 55.8 63.9 67.7 77.1 62.3 68.4 71.9 60.7 71.9 62.6 69.4 67.1 82.6 CLIP-L/14@224px 70.6 64.2 56.1 67.7 73.2 79.0 68.1 71.0 76.1 66.1 74.5 69.4 71.0 69.7 81.9 Finetuned Models VisualBERT 67.1 58.7 59.4 66.5 68.4 77.1 61.3 66.1 71.9 60.7 73.6 63.2 66.5 63.9 82.3 VL-T5 65.8 56.5 61.3 62.9 69.4 77.1 60.3 66.8 70.3 55.8 74.2 61.0 61.0 61.9 82.6 LLaVA-OneVision-7B 67.8 60.0 61.3 64.8 71.3 80.7 61.3 64.8 73.9 62.3 73.6 62.6 67.4 63.9 81.3 CaSigLIP2 (ours) 69.8 62.6 58.7 67.1 72.3 77.7 65.8 72.6 75.8 64.2 75.5 64.2 70.7 69.4 82.6 CaCLIP (ours) 71.5 65.8 60.0 69.4 73.6 81.0 68.7 74.2 75.2 66.8 75.2 66.8 71.0 70.0 83.2 $cIC_{RegionScore\uparrow}$ W/O RAG 40.2 31.3 36.3 36.3 48.8 47.5 27.5 37.5 22.5 67.5 62.5 10.0 35.0 18.8 47.5 SigLIP2-SO/14@384px 52.5 50.0 56.3 36.3 57.5 53.8 58.8 56.3 61.3 70.0 66.3 30.0 53.8 41.3 53.8 CLIP-L/14@224px 55.4 56.3 63.8 41.3 66.3 60.0 61.3 62.5 67.5 70.0 70.0 31.3 55.0 47.5 58.8 Finetuned Models VisualBERT 56.3 58.8 63.8 41.3 58.8 62.5 61.3 61.3 66.3 71.3 70.0 30.0 58.8 45.0 62.5 VL-T5 55.4 60.0 65.0 41.3 62.5 66.3 62.5 58.8 67.5 71.3 68.8 30.0 61.3 40.0 61.3 LLaVA-OneVision-7B 56.0 60.0 65.0 40.0 60.0 62.5 61.3 62.5 66.3 71.3 71.3 28.8 57.5 47.5 60.0 CaSigLIP2 (ours) 56.3 57.5 65.0 36.3 61.3 63.8 58.8 61.3 67.5 71.3 70.0 27.5 57.5 46.3 65.0 CaCLIP (ours) 58.1 60.0 67.5 42.5 65.0 65.0 62.5 60.0 70.0 76.3 75.0 27.5 62.5 52.5 55.0
6 Analysis and Further Discussion
Culture-aware retrieval augmentation substantially benefits VLMs across both cVQA and cIC tasks, compared to their no-RAG counterparts. In this section, we explore the margin of the improvement, cultural preference and the effectiveness of cultural annotation.
6.1 Scaling Models Yields Diminishing or Negative Returns across Retrievers
Figure 5: Performance improvements for smallest and largest models per family with multimodal retrievers. Scaling models yields marginal gains with various retrievers, even negative effects in both cVQA and cIC tasks. "ACC." denotes accuracy; "R.S." refers to the RegionScore; "$\Delta$" represents the change incorporated with RAG compared to the non-RAG baseline.
In the cVQA task, within the same VLM family, performance differences between RAG and non-RAG approaches exhibit non-monotonic trends as model size scales. For all four model families, larger models show marginal or even negative returns from RAG integration. What’s more, sensitivity to RAG varies across model families. Notably, DeepSeek-VL2 demonstrates the most pronounced performance gap: the smallest model benefits from RAG with an average improvement of approximately +5%, whereas the largest model in the same family suffers a degradation of around -6% on average. In the cIC task, the effectiveness of RAG exhibits a consistent trend with respect to model scale within a given model family. Across all four model families evaluated, larger models tend to benefit less—or at most comparably—from the integration of RAG, suggesting diminishing returns at higher capacity. Among them, Gemma3 models show the smallest relative improvement, achieving approximately a +7% performance gain on average, whereas InternVL3 models yield the highest benefit, with performance gains reaching up to +30%.
6.2 Differences across Countries
We evaluate all models using CaCLIP across a range of countries for both tasks, as shown in Figure 6. In the cVQA setting, most VLMs exhibit substantially diminished performance on culture-specific questions regarding Nigeria and Indonesia, in contrast to their performance on questions under other national contexts. Interestingly, questions related to Spanish culture reveal high inter-model variance,
Figure 6: Performance of 14 VLMs with CaCLIP across different countries. Despite the integration of CaCLIP, disparities in model performance persist across countries.
with accuracy differentials reaching up to 35%, underscoring significant discrepancies in cultural representation across models. In the cIC task, VLMs consistently underperform on images and documents associated with Indian cultural contexts, while achieving the highest RegionScores on Korean culture-related inputs. Model performance on Indian culture is particularly volatile, indicating inconsistent cultural grounding across architectures. By comparison, Korean and Chinese cultural inputs yield more stable performance across models, suggesting entrenched model-specific preferences in cultural alignment (see more results in Appendix J).
6.3 Ablation Study on Annotation Questions
Figure 7: Ablation for different annotation questions. Combining all three culture-relevant annotation questions yields the best performance.
We further perform ablation studies across diverse combinations of annotation questions to assess their impact on downstream performance. Specifically, we evaluate 13 open-weight VLMs equipped with either CaSigLIP or CaCLIP, each trained on datasets constructed using varying subsets of culture-relevant annotations. From Figure 7, we can observe that leveraging all three questions (Q1 regarding country association; Q2 for topic alignment; Q3 for visual representation fidelity) yields the strongest performance on both cVQA and cIC tasks. For the cVQA task, we find Q1 provides the most significant benefit to CaSigLIP, whereas CaCLIP gains more from Q2. Among all pairwise combinations, the joint supervision from Q1 (country associations) and Q2 (topic alignment) proves slightly more effective than other pairs. In the cIC task, both CaSigLIP and CaCLIP achieve better performance improvements when trained with data derived from Q1, compared to other single-question sets. For pairwise combinations, CaCLIP benefits most from the Q1+Q3 combination, while CaSigLIP shows a clear preference for the Q2+Q3 setup.
7 Conclusion
We introduce Ravenea, a novel benchmark dataset designed to comprehensively evaluate the cultural sensitivity of diverse multimodal retrievers across 14 SOTA VLMs. The benchmark consists of culturally contextualized queries and image captions from eight countries, curated and paired with Wikipedia passages ranked by human annotators based on their cultural relevance. Our findings highlight the potential of RAG for visual cultural understanding, particularly when lightweight VLMs are enhanced with culturally-aware multimodal retrievers such as CaCLIP, which consistently outperform their non-augmented counterparts – with the caveat that the largest models in each VLM family often exhibit diminishing returns when integrated with RAG. Notably, the inclusion of culture-sensitive questions during data annotation significantly improves effectiveness of multimodal retrievers and enhances performance in downstream tasks.
8 Acknowledgment
We would like to thank all the annotators for their work. Special thanks to Nico Lang, Peter Ebert Christensen, Zhaochong An, Stella Frank, Srishti Yadav, for providing helpful research advice. Serge Belongie and Jiaang Li are supported by the Pioneer Centre for AI, DNRF grant number P1. Ivan Vulić is supported by a personal Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137). Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145).
Appendix A Limitations
While Ravenea establishes a solid foundation for advancing the study of visual culture understanding with retrieval augmentation, it has three limitations that warrant future attention. First, due to budgetary constraints, the dataset’s cultural scope is currently limited to eight countries and eleven categories. Although this selection introduces meaningful diversity, it does not comprehensively represent the global spectrum of cultural perspectives, particularly those of underrepresented or marginalized communities. Second, our use of Wikipedia as the primary external knowledge source introduces inherent biases and may lack the depth, plurality, and contextual richness necessary for nuanced cultural interpretation. Finally, due to resource limitations, we were unable to include certain proprietary VLMs that require paid APIs, such as Gemini 2.5 Pro [52] and Claude Opus 3.7 [53]. We hypothesize that their performance would be comparable to GPT-4.1, which was included in our evaluation, but this remains an open empirical question.
Appendix B Future Directions
Our work opens several avenues for advancing visual culture understanding in multimodal models. First, expanding Ravenea to include more countries, cultural categories, and diverse knowledge sources, beyond Wikipedia, would improve coverage and reduce institutional bias. Second, future benchmarks could include richer tasks beyond cVQA and cIC, such as culture-grounded object recognition, historical retrieval, and symbolic interpretation, to better capture cultural semantics. Third, our results suggest a need for culturally-aware evaluation metrics, particularly for text generation. The limited effectiveness of retrieval augmentation in larger models also warrants further study, especially regarding how cultural knowledge is integrated and utilized. Together, these directions aim to support the development of more culturally-sensitive and globally robust vision-language models.
Appendix C Ethics Statement
This work focuses on improving the cultural awareness of VLMs through retrieval-augmented methods. All data used in the construction of the Ravenea benchmark were sourced from publicly available datasets and Wikipedia, a community-curated open-access knowledge base. To protect individual privacy, we apply automated face detection and blur all identifiable faces in images prior to release. To mitigate cultural bias and ensure broad representation, the benchmark includes images and documents spanning eight countries and eleven cultural domains, curated and annotated by a diverse group of annotators. We provide detailed documentation of the annotation process and guidelines to support transparency and reproducibility. While enhancing cultural understanding is a central goal of this work, we acknowledge that culture is inherently complex, dynamic, and context-dependent. Consequently, the benchmark cannot capture the full richness of any cultural context. Finally, this work does not involve any personally identifiable information (PII), biometric data, or sensitive attributes. All human annotators were compensated fairly, and data annotation adhered to ethical guidelines for responsible research.
Appendix D Experimental Details for Multi-modal Retrievers
Dataset.
We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance. To ensure fair evaluation across regions, we adopt the train-validation-test split strategy from class-imbalance loss [54], yielding an approximate 85-5-10 data split. This setup guarantees that validation and test sets contain a balanced set of images per country (Figure 8), thereby mitigating evaluation bias from skewed geographic distributions. During training, we employ random cropping for images as a sample-level augmentation to normalize the distribution of training instances across countries, further mitigating region-specific sampling biases.
Figure 8: Data distributions across eight countries. Restrict the dataset to the images paired with at least one document annotated by human raters as culturally relevant.
Table 5: Hyperparameters for finetuning on five models.
Hyperparameters VisualBERT VL-T5 LLaVA-OneVision-7B CLIP-L/14@224px SigLIP2-SO/14@384px
batch size 32 128 4 64 64
lr 2e-4 1e-4 1e-4 1e-5 1e-5
lr warmup ratio - 0.1 - - -
weight decay - - - - -
Max Epoch 100 20 10 50 50
Patience 10 10 3 5 5
early stopping Yes
optimizer AdamW
Using LORA - - Yes - -
Hyperparameters.
In this work, we adopt different sets of hyperparameters as VisualBERT, VL-T5, LLaVA-OneVision-7B, CLIP-L/14@224px, SigLIP2-SO/14@384px. For VisualBERT, VL-T5, and LLaVA-OneVision-7B we follow the setting in [44, 45, 46]. We show the training hyperparameters in rerankering experiments for all models in Table 5. All experiments are conducted using a maximum of 2 Nvidia H100 GPUs.
Appendix E Data Statistics
Table 6: Statistics of images in each country.
India Mexico Indonesia China Nigeria Korea Spain Russia
408 349 309 303 223 151 142 77
Table 7: Statistics of images in each category.
Architecture Cuisine History Art Daily Life Companies Sports & Recreation Transportation Religion Nature Tools
403 402 278 275 185 80 73 68 52 31 21
We then apply cryptographic hashing (SHA-256) to identify and remove duplicate images, resulting in a cleaner and more distinct cultural image set. Consequently, Ravenea comprises images collected from eight countries across four continents, spanning eleven distinct categories. The distribution of images by country and by category is detailed in Tables 6 and 7, respectively.
Appendix F Evaluation of VLMs
For closed-weight model, GPT-4.1, we directly call the corresponding API. For open-source models, we use vllm [55]. During the evaluation, to ensure the stability of the results, we set the temperature parameter to 0.0 and the maximum output length to 256. All open-weight models are listed in Table 8. To avoid the impact of the length the retrieved content, we use the first 256 words in the top-1 Wikipedia document.
Table 8: Model details: Hugging Face model names.
Model Hugging Face Model Name
LLaVA-OneVision-7B [19] llava-hf/llava-onevision-qwen2-7b-ov-hf
Phi-4-Multimodal [21] microsoft/Phi-4-multimodal-instruct
Pixtral [20] mistral-community/pixtral-12b
Qwen2.5VL family [23]
Qwen/Qwen2.5-VL-3B-Instruct
Qwen/Qwen2.5-VL-7B-Instruct
Qwen/Qwen2.5-VL-72B-Instruct-AWQ
DeepSeek-VL2 family [25]
deepseek-ai/deepseek-vl2
deepseek-ai/deepseek-vl2-tiny
InternVL3 family [24]
OpenGVLab/InternVL3-2B
OpenGVLab/InternVL3-8B
OpenGVLab/InternVL3-38B
Gemma3 family [22]
google/gemma-3-4b-it
google/gemma-3-27b-it
Appendix G Annotation Details
Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation. We found that directly asking annotators to rate overall cultural relevance on a continuous scale (e.g., 0–10) led to unreliable and inconsistent labels. This difficulty arises from several factors: (1) the semantic meaning of intermediate scores is ambiguous, (2) annotators tended to overemphasize a few salient visual elements [56], and (3) small numerical differences (e.g., between 5 and 6) often fail to reflect meaningful distinctions, especially given the cognitive load of processing lengthy Wikipedia documents, resulting in intra-annotator variance even on repeated examples. Instead, given an image–caption–document triplet, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: country association, topic alignment, and explicit visual representation.
Prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines. To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks. For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25. They are asked to assess whether each article provides meaningful background or contextual information that is directly relevant to the cultural elements described in the caption or depicted in the image.
Appendix H A New Metric: RegionScore
To evaluate the extent to which captions reference specific geopolitical regions, we introduce a RegionScore. This metric measures whether a caption contains explicit references to a country or its common demonyms. $N$ be the total number of samples. $\text{PRED}_{i}$ denote predicted captions for the $i$-th sample. $C_{i}$ be the country name associated with the $i$-th sample. $\text{Adj}(C_{i})$ denote the set of adjectives or demonyms associated with $C_{i}$. $\mathcal{T}_{i}=\{C_{i}\}\cup\text{Adj}(C_{i})$ denote the set of region-related terms for the $i$-th sample. We define binary indicators for each sample:
$\delta^{\text{pred}}_{i}=\begin{cases}1&\text{if any }t\in\mathcal{T}_{i}\text{ appears in }\text{PRED}_{i}\\ 0&\text{otherwise}\end{cases}$ (5)
The RegionScore for predicted captions are then computed as:
$\text{RegionScore}_{\text{PRED}}=\frac{1}{N}\sum_{i=1}^{N}\delta^{\text{pred}}_{i}$ (6)
These scores reflect the proportion of captions that include explicit regional identifiers. A higher score indicates stronger region-awareness in the captioning. For the ground truth in CCUB dataset, the $\text{RegionScore}_{\text{GT}}=99\%$.
Appendix I Details of Human Evaluation in the cIC Task
We randomly sample 10 images and generate captions using 14 vision-language models (VLMs) under three configurations: CaCLIP, CLIP, and a no-retrieval baseline, yielding 420 captions in total. Four expert annotators participated in the evaluation. For each image, they were presented with caption triplets—one per retrieval configuration—and asked to assess approximately 35 such triplets each. Annotators are instructed to select the caption that most accurately and appropriately reflects the cultural context depicted. To evaluate annotation consistency, we randomly sample 30 triplets and assign them to a fifth expert annotator. Inter-annotator agreement (IAA) between this annotator and the original four, measured using Cohen’s $\kappa$, is 0.595.
Appendix J Additional Results
Figure 9: Performance of the 14 VLMs equipped with CaCLIP in cVQA task.
As illustrated in Figures 9 and 10, the models exhibit diverse cultural preferences. Notably, most models achieve relatively stronger performance on Chinese and Indian cultural contexts in the cVQA task, and on Chinese and Korean contexts in the cIC task.
Figure 10: Performance of the 14 VLMs equipped with CaCLIP in cIC task.
We report the CIDEr, ROUGE-L, BERTScore, and CLIPScore metrics for 14 VLMs on the cIC task in Table 9.
Table 9: Four metrics comparison of RAG and Non-RAG methods for CCUB task.
Open Weights Closed Weights Method Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $CCUB_{Rouge-L}$ W/O RAG 18.3 18.7 18.2 21.8 18.2 16.5 17.8 19.6 19.8 16.7 15.9 20.6 15.4 21.4 19.6 Frozen Models SigLIP2-SO/14@384px 18.0 20.0 19.1 23.8 17.2 14.9 19.2 17.4 17.5 15.1 14.9 23.1 14.0 23.6 17.0 CLIP-L/14@224px 18.1 20.9 18.5 23.0 16.8 14.5 18.3 17.2 17.5 15.1 15.8 22.4 14.4 24.7 16.1 Finetuned Models VisualBERT 18.6 19.5 18.3 22.5 16.8 14.8 17.4 17.4 16.9 15.0 16.0 22.5 14.8 24.3 19.0 VL-T5 17.6 19.0 18.1 23.1 17.9 14.4 18.2 17.2 16.9 14.7 15.4 20.8 13.8 23.5 17.4 LLaVA-OneVision-7B 17.3 18.4 17.7 23.0 17.2 14.5 17.6 16.4 17.1 14.6 15.0 20.9 13.1 22.8 17.3 SigLIP2-SO/14@384px 18.4 19.3 19.2 24.5 18.5 14.8 18.4 18.2 17.2 15.2 16.4 22.2 14.3 23.5 18.9 CLIP-L/14@224px 18.1 19.6 18.6 23.9 17.1 14.7 18.0 17.3 17.7 16.0 16.3 21.5 13.7 24.2 17.0 $CCUB_{CIDER}$ W/O RAG 28.1 30.0 27.2 49.2 22.9 16.9 10.5 18.6 35.8 11.2 10.1 49.1 16.7 47.4 34.4 Frozen Models SigLIP2-SO/14@384px 25.2 39.2 20.5 58.5 16.9 7.6 12.4 7.1 6.3 4.7 6.1 62.0 10.1 51.4 17.2 CLIP-L/14@224px 23.8 40.9 18.8 59.6 13.0 6.2 9.6 7.8 4.8 4.7 6.6 56.3 11.1 49.7 14.6 Finetuned Models VisualBERT 26.1 40.2 24.0 60.9 16.3 8.2 10.0 7.5 5.1 6.4 10.2 51.1 15.1 49.5 24.4 VL-T5 24.5 37.0 20.6 51.1 16.0 7.3 10.2 5.0 4.0 6.8 7.5 52.5 10.3 48.4 21.6 LLaVA-OneVision-7B 21.7 38.8 18.5 62.6 13.2 4.8 8.4 5.7 3.4 5.7 4.5 48.9 11.4 42.2 17.6 SigLIP2-SO/14@384px 28.2 40.2 22.8 64.7 17.8 8.4 9.7 5.8 4.4 4.7 7.7 62.3 13.2 54.8 26.5 CLIP-L/14@224px 24.6 35.4 22.1 61.5 15.7 8.9 9.9 6.5 5.8 5.8 7.8 53.1 11.2 51.7 18.3 $CCUB_{BERTScore}$ W/O RAG 54.7 55.4 55.6 55.6 55.7 54.2 52.5 54.5 55.4 54.1 54.8 53.5 52.7 56.6 56.8 Frozen Models SigLIP2-SO/14@384px 54.9 57.2 56.4 57.9 55.2 53.0 55.1 54.8 55.0 52.4 54.0 56.9 52.6 58.7 54.8 CLIP-L/14@224px 54.8 57.6 56.2 57.7 54.8 52.4 54.8 55.1 55.3 52.6 53.7 55.9 52.5 59.4 54.6 Finetuned Models VisualBERT 54.8 57.1 56.5 57.6 54.9 52.8 54.6 54.9 54.4 52.1 53.8 55.7 52.8 58.5 56.2 VL-T5 54.2 56.7 56.0 58.1 54.9 52.6 53.9 54.0 54.4 52.0 53.4 54.4 52.1 57.8 55.7 LLaVA-OneVision-7B 54.2 56.6 55.5 58.2 54.5 52.1 53.9 54.2 54.6 51.6 53.3 55.0 51.8 58.3 54.9 SigLIP2-SO/14@384px 55.2 57.4 57.1 57.9 55.7 53.2 54.7 55.2 54.9 52.5 54.2 56.8 52.6 58.7 56.7 CLIP-L/14@224px 55.0 57.0 56.5 58.2 54.8 53.2 54.6 55.3 55.3 52.9 54.2 56.0 52.3 59.3 55.4 $CCUB_{CLIPScore}$ W/O RAG 19.1 19.1 18.8 19.6 18.3 18.2 18.4 18.7 19.2 18.4 18.4 20.2 18.5 20.2 18.7 Frozen Models SigLIP2-SO/14@384px 19.1 19.3 19.0 19.9 18.9 18.8 19.1 18.8 18.7 18.6 18.5 20.5 18.8 20.0 18.6 CLIP-L/14@224px 19.0 19.2 19.0 20.0 19.0 18.5 19.2 18.8 18.6 18.8 18.4 20.4 18.5 19.6 18.3 Finetuned Models VisualBERT 19.1 19.1 18.9 20.2 18.9 18.9 19.1 18.9 18.8 18.5 18.6 20.3 18.7 19.9 18.5 VL-T5 19.1 19.4 19.1 19.9 19.0 18.8 19.0 19.1 18.7 18.5 18.7 20.4 18.7 19.9 18.4 LLaVA-OneVision-7B 19.1 19.3 18.9 20.1 18.9 18.7 19.3 18.9 18.6 18.5 18.6 20.4 18.8 19.7 18.6 SigLIP2-SO/14@384px 19.1 19.2 18.9 20.1 18.8 18.6 19.1 19.0 18.5 18.6 18.5 20.4 18.8 19.8 18.5 CLIP-L/14@224px 19.0 19.2 18.9 20.0 19.1 18.6 19.2 18.9 18.7 18.6 18.4 20.1 18.8 19.7 18.6
Figure 11: Average improvement across 14 VLMs in different countries with 4 retrievers for cVQA task.
As shown in Figures 12 and 11, VLMs demonstrate varying degrees of performance shifts across cultural contexts in both the cIC and cVQA tasks. In the cIC setting, most fine-tuned retrievers yield noticeable improvements over their original counterparts, indicating the effectiveness of retrieval adaptation. Performance on the cVQA task reveals more nuanced outcomes. While fine-tuned retrievers generally exhibit large performance improvement compared to non-RAG baselines, certain countries, such as Spain and Indonesia, experience exhibit diminishing returns. These discrepancies may stem from the limited presence of culturally representative visual content in the training data. Although targeted image augmentation strategies were employed to alleviate this imbalance, the results suggest that data distribution remains a significant bottleneck. Understanding and addressing such cross-cultural performance disparities in multimodal tasks like cVQA will be an important direction for future work.
Figure 12: Average improvement across 14 VLMs in different countries with 4 retrievers in cIC task.
Appendix K Human Annotation & Evaluation Interface
The annotation interface are shown in Figure 14 and 15. The interface for human evaluation in cIC task is shown in Figure 13.
Figure 13: Human evaluation interface for cIC task.
Figure 14: Human annotation instructions.
(a) Annotation interface (1).
(b) Annotation interface (2).
Figure 15: Human annotation interfaces.
Appendix L Correlation between Automatic Metrics and Human Judgments
To assess the correlation between human preferences and automatic evaluation metrics, we compute Kendall’s Tau rank correlation. Human annotations are segmented according to the output chunks produced by corresponding 14 VLMs with each retriever variant (CaCLIP-based, CLIP-based, and non-RAG). Within each segment, we calculate the selection winning ratio for each retrieval method, yielding a human preference vector formed by concatenating these ratios across all evaluation instances. For the automatic evaluation, we extract BERTScore, CIDEr, ROUGE-L, and CLIPScore for each corresponding retrieval variant and VLM. These scores are similarly concatenated into a metric-based vector. Finally, we compute Kendall’s Tau between the human and metric vectors to quantify the consistency between automatic rankings and human judgments.
Appendix M Prompts
The prompts that we used to collect the documents and two downstream tasks were designed to ensure consistency across different models. One prompt example for cultural caption generation is shown in Table 10. The prompts for both with and without retrieval augmentation for VLMs are shown in Table 12 and 11.
Table 10: Prompt example for generating captions used to retrieve the Wiki documents.
Prompt for data collection GPT-4o SYSTEM: Default
USER:
Generate a culture related caption given the image with around 2 sentences. Please include the name of the thing shown in the picture within the caption if it’s strongly related to the culture. Name the three most culturally relevant entities in the image and attach the name at the end. Follow the format of: caption. entity name1; entity name2; entity name3.
Table 11: Prompt examples without RAG. Multimodal prompt samples with interleaved image are shown for both CVQA and CCUB tasks.
CVQA task CCUB task
SYSTEM:
You are a helpful assistant.
USER:
Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
Question: Which city is famous for such artwork?
A) Zhejiang Yiwu
B) Jingdezhen, Jiangxi
C) Datong, Shanxi
D) Zibo, Shandong
SYSTEM:
You are a helpful assistant.
USER:
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content and culturally relevant elements of the image. Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph..."). Do not generate multiple options.
Table 12: Prompt examples with RAG. Multimodal prompt samples with interleaved image are shown for both CVQA and CCUB tasks.
CVQA task CCUB task
SYSTEM:
You are a helpful assistant.
USER:
Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
The scope of the question is strictly limited to the given image. However, please analyze and incorporate information from both the image and the following document to answer the question.
Document:
Buildings and structures Buildings about 800 – Borobudur temple in Java completed. 802 Haeinsa of Korea, is constructed. Palace of Charlemagne in Aachen, Carolingian Empire completed (begun about 790). The Palatine Chapel still stands. At Oviedo in the Kingdom of Asturias Cámara Santa constructed. First reconstruction of Oviedo Cathedral begun by Tioda. 815 – Second Temple of Somnath built in the Pratihara Empire, India. 816 – Reims Cathedral begun. 810s – Chapel of San Zeno in Santa Prassede, Rome decorated. 818 – Old Cologne Cathedral built....
Question: What is the name of the square where the cathedral and the statue of the image are located?
A) Riego Square
B) Trascorrales Square
C) The square of Alfonso II the Chaste
D) The Fontan square
SYSTEM:
You are a helpful assistant.
USER:
Write a concise, one-sentence caption for the given image. The generated caption must contain the visual content and culturally relevant elements of the image. Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph..."). Do not generate multiple options. Please consider the following context:
Architecture of Nigeria was historically influenced by environmental conditions as well as social and cultural factors. The coming of missionaries and political changes brought about by colonialism precipitated a change in architectural style and utility of buildings. A Gothic revival style was adopted for early churches built in the colony of Lagos. A one or two storey timber house building made with pre-fabricated material components and designed with the influence of classic antiquity styles served as mission house for the missionaries...

### v2

@@
 Prior work has demonstrated the effectiveness of retrieval-augmented generation (RAG) in enhancing cultural understanding in text-only settings, while its application in multimodal scenarios remains underexplored.
-To bridge this gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a new benchmark designed to advance visual culture understanding through retrieval, focusing on two tasks: culture-focused visual question answering (cVQA) and culture-informed image captioning (cIC).
-Ravenea extends existing datasets by integrating over 10,000 Wikipedia documents curated and ranked by human annotators.
-With Ravenea, we train and evaluate seven multimodal retrievers for each image query, and measure the downstream impact of retrieval-augmented inputs across fourteen state-of-the-art VLMs.
-Our results show that lightweight VLMs, when augmented with culture-aware retrieval, outperform their non-augmented counterparts (by at least $3.2\%$ absolute on cVQA and $6.2\%$ absolute on cIC).
-This highlights the value of retrieval-augmented methods and culturally inclusive benchmarks for multimodal understanding.
+To bridge this gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a new benchmark designed to advance visual culture understanding through retrieval, focusing on two tasks: culture-centric visual question answering (cVQA) and culture-informed image captioning (cIC).
+Ravenea extends existing datasets by integrating 11,396 unique Wikipedia documents curated and ranked by human annotators.
+Through the extensive evaluation on seven multimodal retrievers and seventeen VLMs, Ravenea reveals some undiscovered findings: (i) In general, cultural grounding annotations can enhance multimodal retrieval and corresponding downstream tasks.
+(ii) VLMs, when augmented with culture-aware retrieval, generally outperform their non-augmented counterparts (by averaging $6\%$ on cVQA and $11\%$ on cIC).
+(iii) Performance of culture-aware retrieval augmented varies widely across countries.
+These findings highlight the limitations of current multimodal retrievers and VLMs, underscoring the need to enhance visual culture understanding within RAG systems.
+We believe Ravenea offers a valuable resource for advancing research on retrieval-augmented visual culture understanding.
+Project page: jiaangli.github.io/ravenea
 1 Introduction
-Vision-language models (VLMs) are increasingly deployed in real-world applications, from education to assistive technologies di2024educational ; de2024vqask ; bates-etal-2025-template ; yao2025countllm , where understanding not only visual content but also the surrounding cultural context is crucial.
-Despite achieving impressive performance on general tasks yue2024mmmu ; lin2014microsoft ; an2025multimodality ; an2025generalized , VLMs often struggle to capture cultural nuances, such as traditions, symbols, and region-specific practices that require external, culturally grounded knowledge (nayak-etal-2024-benchmarking, ; liu2025culturevlm, ; khanuja-etal-2024-image, ; cao2024exploring, ).
-For example, as shown in Figure 1, a VLM may incorrectly identify the season of a festival scene as ‘Autumn’, overlooking that the image depicts Kyoto’s Gion Festival, which occurs in July and corresponds to ‘Summer’.
-A promising approach to address this limitation is the integration of external knowledge through retrieval-augmented generation (RAG) lewis2020retrieval , which has shown success in improving cultural awareness in language models (seo2025valuesrag, ; lertvittayakumjorn2025towards, ).
-However, prior work in the culture domain has predominantly been confined to text-only settings.
-Meanwhile, existing culture-related multimodal datasets primarily focus on evaluating VLM outputs on culturally oriented tasks, with limited emphasis on the integration of external cultural knowledge.
-As a result, the potential of RAG to improve multimodal cultural understanding remains underexplored.
-Figure 1: Effectiveness of culture-aware RAG.
-Given a culturally grounded visual question, VLMs enhanced with culture-aware RAG—retrieving relevant Wikipedia documents—generate more accurate answers than their non-RAG counterparts.
-(see Section 5.2).
-To bridge the gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a manually curated dataset designed to evaluate cultural understanding in VLMs with retrieval support.
-We construct Ravenea based on two existing datasets: CVQA (romero2025cvqa, ), which includes culturally relevant visual questions and corresponding answers, and CCUB (liu2023towards, ), offering culturally contextualized captions to foster inclusivity in text-to-image generation  [^1] We reuse the cultural captions from CCUB as ground-truth references for the inverse task, image-to-text generation, specifically for culture-aware image captioning..
+Vision-language models (VLMs) are increasingly deployed in real-world applications, from education to assistive technologies (Baker et al., 2021; Di Nuovo et al., 2024; De Marsico et al., 2024; Karamolegkou et al., 2025), where understanding not only visual content but also the surrounding cultural context is crucial.
+Despite achieving impressive performance on general tasks (Lin et al., 2014; Yue et al., 2024; Bai et al., 2025a), VLMs often struggle to capture cultural nuances, such as traditions, symbols, and region-specific practices that require external, culturally grounded knowledge (Nayak et al., 2024; Khanuja et al., 2024; Liu et al., 2025; Winata et al., 2025).
+For example, when analyzing images of cultural events, VLMs may misinterpret region-specific attire as generic clothing, thereby overlooking its ceremonial significance (Nayak et al., 2024).
+What’s more, they also risk reinforcing existing cultural biases, as biased collections or metadata frequently lead VLMs to perform much better in mainstream culture contexts (e.g.
+western culture) while marginalizing minority traditions (Nayak et al., 2024).
+A promising approach to mitigate the limitations is the integration of external knowledge through retrieval-augmented generation (RAG) (Lewis et al., 2020), which has shown success in improving cultural awareness in language models (Seo et al., 2025; Lertvittayakumjorn et al., 2025).
+However, prior work on culture has largely remained in text-only settings, while existing multimodal culture-related datasets mostly probe VLMs’ memorized cultural knowledge instead of testing their understanding of culture in real-world contexts.
+Consequently, it remains unclear to what extent current mutlimodal retrievers support reliable cultural retrieval, and whether RAG can effectively advance multimodal cultural understanding for VLMs.
+This shortfall underscores the need for a dedicated benchmark that probes various facets of visual culture comprehension for multimodal retrievers and VLMs.
+Figure 1: Samples from Ravenea.
+Ravenea evaluates the multimodal culture-awareness RAG with VLMs across eight countries spanning four continents: Asia, Africa, Europe, and Latin America.
+It comprises two tasks designed to assess visual culture understanding abilities and the effect of extra cultural knowledge across different countries.
+With 1,868 instances and 11,369 human-ranked documents, Ravenea covers eleven real-world scenarios, including architecture, cuisine, art, etc.
+To bridge the gap, we introduce Ravenea (Retrieval-Augmented Visual culturE uNdErstAnding), a manually curated dataset designed to evaluate multicultural understanding in VLMs with retrieval support.
+We construct Ravenea based on two existing datasets: CVQA (Romero et al., 2025), which includes culturally relevant visual questions and corresponding answers, and CCUB (Liu et al., 2023), offering culturally contextualized captions to foster inclusivity in text-to-image generation.
+[^1] We reuse the cultural captions from CCUB as ground-truth references for the inverse task, image-to-text generation, specifically for culture-aware image captioning.
 For each instance drawn from the source datasets, we append a set of Wikipedia documents that have been human-ranked based on their cultural relevance to the associated image.
-This curation effort, designed to ensure broad cultural representation, contains data related to eight countries and spans eleven diverse categories, comprising more than 1,800 images and 10,000 human-ranked documents.
-Ravenea thus provides a retrieval-augmented benchmark for evaluating cultural sensitivity in multimodal retrievers, and further allows for assessing how well VLMs integrate and make use of retrieved cultural context.
-Specifically, we focus on two culturally grounded tasks: (i) culture-focused visual question answering (cVQA) and (ii) culture-informed image captioning (cIC).
-We select these two tasks for their centrality in assessing cultural understanding in VLMs—question answering tests context-aware reasoning, while captioning evaluates generation sensitivity to cultural nuances.
-With Ravenea, we first train and evaluate seven multimodal retrievers that use both visual and textual inputs to retrieve Wikipedia documents for a given query image based on their cultural relevance.
-Then, we evaluate a diverse set of state-of-the-art (SOTA) VLMs, including GPT-4.1 achiam2023gpt , LLaVA-OneVision-7B li2025llavaonevision , Pixtral agrawal2024pixtral , Phi-4 Multimodal abouelenin2025phi , the Gemma3 family team2025gemma , the Qwen2.5-VL family bai2025qwen2 , the InternVL3 family zhu2025internvl3 , and the Deepseek-VL2 family wu2024deepseek , each with and without multimodal retrieval, to assess the impact of retrieval augmentation on cultural understanding.
-Our dataset provides a testbed for assessing the cultural relevance capabilities of multimodal retrievers and the effectiveness of VLMs in consuming and using such retrieved cultural context.
+This curation effort, designed to ensure broad cultural representation, contains data related to eight countries and spans eleven diverse categories (e.g.
+architecture, art, cuisine, etc), comprising 18,680 human-labeled image-document pairs.
+Ravenea thus provides a multimodal retrieval-augmented benchmark for evaluating multicultural sensitivity of various retrievers, and further allows for assessing how well VLMs utilize retrieved cultural context in their reasoning process.
+Specifically, as shown in Figure 1, we focus on two culturally grounded tasks: (i) culture-centric visual question answering (cVQA) and (ii) culture-informed image captioning (cIC).
+With Ravenea, we first train and evaluate seven multimodal retrievers that use both visual and textual inputs to retrieve Wikipedia documents for a given image query based on their cultural relevance.
+We then evaluate 17 widely used VLMs, both with and without multimodal retrieval, spanning open-source and proprietary models as well as sizes from 2B to 78B parameters, to assess the effect of retrieval augmentation on cultural understanding in the cVQA and cIC tasks.
+Our benchmark provides a testbed for evaluating the cultural relevance capabilities of multimodal retrievers and the effectiveness of VLMs in consuming and using such retrieved cultural context.
 Our contributions and key findings include:
-  • Ravenea benchmark: We introduce Ravenea, the first benchmark aimed at evaluating VLMs and multimodal retrieval in leveraging external knowledge for visual culture understanding.
-It comprises a diverse, large-scale collection of human-curated, culturally related documents linked to images from eight countries across eleven categories, enabling evaluation on two tasks: culture-focused visual question answering and culture-informed image captioning.
-(Section 3)
+  • Ravenea benchmark: We introduce Ravenea, the first benchmark for evaluating VLMs in using external knowledge for visual culture understanding.
+It covers eight countries across eleven categories, linking images to human-ranked Wikipedia documents on two tasks: cVQA and cIC.
   • Cultural grounding annotations enhance multi-modal retrieval: We evaluate seven retrievers that integrate visual and textual cues to retrieve culturally relevant documents.
 We find fine-tuning retrievers on culture-targeted annotations leads to marked gains in retrieval accuracy, highlighting the value of explicit cultural supervision.
-(Section 6.3)
-  • Benefits of culture-aware retrieval: Culture-aware retrieval boosts task performance across VLMs, with lightweight models showing the greatest improvement.
+  • Benefits of culture-aware retrieval: Culture-aware retrieval boosts task performance across VLMs, with lightweight models showing the better improvement.
 This suggests that such retrieval can seamlessly integrate into downstream VLM tasks, enhancing their performance.
-(Section 5.2)
-  • Cross-cultural variation: Evaluation across eight countries reveals that VLMs exhibit distinct cultural preferences, with each model favoring different regional contexts—suggesting model-specific cultural biases.
-(Section 6.2)
-Figure 2: Ravenea: A Multimodal Retrieval-Augmented Visual culturE uNdErstAnding dataset.
-Left: Examples of cVQA and cIC tasks.
-Middle: Geographic and categorical distribution of cultural references.
-Right: Performance comparison of 14 VLMs, evaluated with and without integration of our culture-aware retriever.
-Here, CaCLIP="culture-aware CLIP-L/14@224px ".
-2 Related Work
-Retrieval augmentation for cultural understanding.
-Retrieval augmentation has demonstrated significant efficacy for culture-related NLP tasks (conia-etal-2024-towards, ; hu-etal-2024-bridging, ; fung-etal-2023-normsage, ).
-Prior work uses sources like the World Values Survey for cultural question answering seo2025valuesrag ; yadav-etal-2025-beyond , or retrieves web and knowledge base content to enhance cultural contextualization lertvittayakumjorn2025towards .
-In multimodal settings, however, cultural retrieval remains underexplored.
-While some approaches retrieve culture-relevant images to fine-tune VLMs liu2025culturevlm , they require additional training.
-In contrast, our method introduces a plug-and-play retrieval system that enriches cultural grounding at inference time without modifying the base model.
-Vision-language culture datasets.
-Several recent studies examine cultural understanding in VLMs through multicultural VQA liu-etal-2021-visually ; nayak-etal-2024-benchmarking , cuisine recognition winata-etal-2025-worldcuisines ; li-etal-2024-foodieqa , and concept-based image retrieval bhatia-etal-2024-local .
-Others benchmark cultural entity recognition using Wikipedia-based prompts nikandrou-etal-2025-crope or curate culture-specific image sets for value-based tasks yadav-etal-2025-beyond .
-While these efforts highlight current VLM limitations, they often rely on static, manually curated datasets.
-A natural extension to address these deficiencies is through culture-relevant multimodal retrieval—an area lacking dedicated datasets.
-Our work specifically aims to fill this critical research gap.
-3 Ravenea Dataset
-We construct our dataset by building culturally relevant document lists for each image in two existing culture-grounded datasets: CVQA romero2025cvqa , a widely used dataset for culture-focused VQA, and CCUB liu2023towards , a dataset designed to mitigate cultural bias in text-to-image generation with culturally contextualized captions.
-To ensure broad geographic and cultural coverage, we curate a diverse subset comprising images from seven countries in CVQA: China, Nigeria, Russia, Spain, Mexico, India, and Indonesia, and all five countries in CCUB: China, Korea, India, Mexico, and Nigeria (see Figure 2).
-After that, we separate the dataset construction process into three critical stages shown in Figure 3.
-3.1 Dataset Construction
-Data collection.
-The data collection process consists of two main steps: (i) culture-related captioning and (ii) document retrieval.
-For culture-related captioning, we generate culturally grounded captions for each image to facilitate more effective attachment of relevant documents.
-Since the CVQA lacks captions and the CCUB provides only brief descriptions, we employ GPT-4o to generate richer, culturally informative captions (see the prompt example in Table 10).
-For document retrieval, we first conduct a coarse filtering using the generated cultural captions as queries for a BM25 Robertson2009ThePR retriever to extract the semantically relevant documents from a large-scale corpus comprising over six million English Wikipedia documents  [^2] https://huggingface.co/datasets/wikimedia/wikipedia.
-To mitigate the impact of inaccurate captions and ensure precise document relevance, we then perform human annotation on the retrieved documents.
+  • Cross-cultural variation: Evaluation across eight countries reveals that VLMs, both with and without cultural RAG exhibit distinct cultural preferences, with each model favoring different regional contexts, suggesting model-specific cultural biases.
+2 Ravenea Construction
+Figure 2: Ravenea: A Multimodal Retrieval-Augmented Visual culturE uNdErstAnding dataset. Geographic and categorical distribution of cultural references in Ravenea.
+We present Ravenea, the first benchmark explicitly designed to assess how retrieval augmentation influences the visual cultural understanding capabilities of VLMs.
+It is derived originally from 30 countries; see Table 4.
+Each image is paired with 10 algorithm-ranked documents based on cultural relevance.
+To ensure high data quality while maximizing linguistic diversity across major language families, we sample data from eight countries spanning multiple continents: China, Nigeria, Russia, Spain, Mexico, India, Indonesia, and Korea.
+[^2] List of languages by total number of speakers.
+We then apply human re-ranking–based annotation to the collected documents to construct Ravenea.
+Consequently, Ravenea comprises 1,868 instances, including 2,331 image-text questions, 655 image-caption pairs, and 18,680 image-document pairs, supporting two tasks: cVQA and cIC, respectively.
+2.1 Taxonomy
+Motivated by the increasing need for VLMs to understand cultural context (Yadav et al., 2025a), we categorize the cultural understanding capability into two core tasks and several detailed axes, as shown in Figure 2.
+Additional dataset statistics are provided in Appendix D.
+cVQA.
+This task evaluates a model’s ability to perceive and reason over culturally grounded visual content, often requiring external, culturally specific knowledge beyond the image itself.
+Building on the CVQA dataset (Romero et al., 2025), which contains culture-related QA pairs, we augment it with culturally relevant documents that supply the additional knowledge required for RAG evaluation.
+Therefore, each instance contains an image, a question, top 10 human-ranked documents based on culture relevance, and multiple-choice options, with only one correct answer.
+cIC.
+This task assesses the VLM’s ability to generate captions that are sensitive to and incorporate cultural nuances.
+It evaluates the model’s generation sensitivity by requiring it to produce culturally contextualized captions for a given image, leveraging information from human-curated, culturally related documents.
+The task reuses cultural captions from the CCUB dataset (Liu et al., 2023) as ground-truth references for this task.
+As a result, for cIC, each instance consists of an image, a human-written caption, and top 10 culturally-related documents.
 Figure 3: Ravenea construction pipeline.
@@
 Right: Postprocessing methods for quality control.
-Relevance annotation.
+2.2 Dataset Curation
+Data collection.
+As shown in Figure 3, the data collection process consists of two main steps: (i) culture-related captioning and (ii) document retrieval.
+For culture-related captioning, we generate culturally grounded captions for each image to facilitate more effective attachment of relevant documents.
+Since the CVQA lacks captions and the CCUB provides only brief descriptions, we employ GPT-4o [^3] Knowledge cutoff: Oct 01, 2023.
+(Achiam et al., 2023) to generate richer, culturally informative captions (see the prompt example in Table 13).
+[^4] We used LLMs to generate culture-informed captions solely as auxiliary references for annotators; these LLM outputs were not directly included in the final dataset.
+For document retrieval, we begin with a coarse-grained filtering stage in which the generated cultural captions serve as queries to a BM25 retriever (Robertson and Zaragoza, 2009).
+This stage retrieves a pool of semantically relevant candidates from a large-scale corpus of over six million English Wikipedia articles.
+[^5] Wikimedia/wikipedia To mitigate the impact of inaccurate captions and ensure precise document relevance, we then perform human annotation on the retrieved documents.
+Cultural relevance annotation.
 Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation.
 For each image-caption pair, annotators are presented with the top 10 Wikipedia documents retrieved by BM25.
-They are asked to assess whether each document provides meaningful background or contextual information that is relevant to the culture described in the caption or the image (see Appendix K).
-Specifically, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: Country association: Is the topic of the Wikipedia article associated with the same country as the image and its caption?
+They are asked to assess whether each document provides background or information that is relevant to the culture described in the caption or the image (see interfaces in Appendix L).
+Specifically, we decomposed cultural relevance into three interpretable and independently verifiable dimensions:
+Country association: Is the topic of the Wikipedia article associated with the same country as the image and its caption?
 Topic alignment: Does the topic of the Wikipedia article align with the semantic category of the image and its caption?
@@
 Additionally, annotators are also instructed to include the title and URL of any relevant Wikipedia article they believe is missing from the top-10 retrieved results.
-These manually suggested articles are treated as the most cultural references closely related to the given image (see details in Appendix G).
+These manually suggested articles are treated as the most culturally relevant references for the given image.
+Detailed annotation statistics, alongside positive, negative and missing ratios by countries, are presented in Appendix E.
 Quality control.
-To ensure the quality and consistency of our annotations, we implement several quality control methods.
-First of all, prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines.
-Table 1: Statistics of the Ravenea dataset.
-The dataset is constructed by curating existing sources and augmenting them with over 10,000 wiki-derived documents to broaden cultural knowledge coverage and enhance content diversity.
-Dataset Images Documents Pairs Questions Captions
-CVQA 1,213 8,319 12,130 2,331 -
-CCUB 655 4,441 6,550 - 655
-Ravenea 1,868 11,580 18,680 2,331 655
-To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks (see Figure 14).
-We also perform an additional quality check on a subset of the dataset.
-Specifically, for each selected countries, we employ an additional local quality checker who is tasked with manually reviewing the annotations to verify their accuracy and adherence to the guidelines.
-The quality checker reviews a random sample of annotated items, focusing on both the relevance labels and the justification behind any edge cases, such as borderline relevance or use of the “Cannot be determined” label.
-If inconsistencies or deviations from the annotation guidelines are identified, the affected samples are flagged for re-annotation.
-The overall acceptance rate from the meta quality checkers is 98.2%.
-The inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$artstein2017inter between the meta checker and annotator on the sampled annotations is 0.83.
-Figure 4: Examples demonstrating the impact of CaCLIP Wikipedia retrieval integration on cVQA and cIC tasks using DeepseekVL2-Tiny.
-When augmented with culture-aware retrieval, the model exhibits enhanced sensitivity to cultural context.
-3.2 Dataset Statistics
-We present the statistics of our dataset in Table 1.
-The dataset comprises a total of 1,868 culturally diverse images, with approximately 65% originating from CVQA and the remaining 35% from CCUB.
-Each image is paired with a GPT-4o-generated cultural caption, as well as the top-10 ranked Wikipedia documents retrieved via a cultural relevance scoring pipeline (Figure 3), yielding a total of 18,680 image-document pairs (see illustrated examples in Figure 4).
-The collection spans eight countries and encompasses a broad spectrum of cultural domains, such as traditional attire, festivals, architecture, cuisine, and social practices (see more statistics in Appendix E).
-4 Culture-aware Multimodal Retriever
-Leveraging the Ravenea dataset, we train and evaluate seven multimodal retrievers to retrieve culturally relevant Wikipedia documents using both visual and textual inputs.
-We fine-tune five representative models—spanning both generative and discriminative paradigms—to optimize multimodal document retrieval.
-Performance is evaluated using standard retrieval metrics, including Mean Reciprocal Rank (MRR) shi2012climf , Precision@k (P@k) jarvelin2017ir , and Normalized Discounted Cumulative Gain (nDCG@k) wang2013theoretical , where $k\in\{1,3,5\}$.
-We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance.
-We fine-tune a VisualBERT-based li2019visualbert ; li-etal-2020-bert-vision reranker following standard BERT-style setups nogueira2019passage , and adapt two multimodal generators—VL-T5 cho2021unifying and LLaVA-OneVision-7B li2025llavaonevision —for end-to-end document retrieval chen2022corpusbrain ; yuan2024asking ; sun2023learning .
-To enhance cultural awareness in the contrastive retrieval, we introduce Culture-Aware Contrastive (CAC) learning, a supervised learning framework compatible with both CLIP and SigLIP architectures.
-We denote the culture-aware fine-tuned versions of CLIP-L/14@224px and SigLIP2-SO/14@384px using CAC as CaCLIP and CaSigLIP2, respectively.
+For caption quality, we employ two independent local annotators assessed both objective accuracy and subjective quality (naturalness and completeness) in a subset, achieving 92% and 94% accuracy with an inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$(Artstein, 2017) of 0.85, and yielding high subjective ratings (naturalness: 4.98/5, completeness: 4.66/5).
+The results demonstrate that the captions are reliable for downstream retrieval of Wikipedia documents.
+For annotation quality, a separate group of annotators were trained with detailed guidelines and a mock test before performing the actual annotation.
+A meta quality checker further validated a random subset, achieving a 98.2% acceptance rate.
+The IAA Cohen’s Kappa ($\kappa$) between the meta checker and annotator on the sampled annotations is 0.83.
+More details are presented in Appendix F.
+3 Experimental Setup
+With Ravenea, we train and evaluate seven multimodal retrievers to retrieve culturally relevant Wikipedia documents using multimodal inputs, spanning both generative and discriminative paradigms.
+To enhance cultural awareness in the contrastive retrieval, we introduce Culture-Aware Contrastive (CAC) learning (details in Section 3.2), a supervised learning framework compatible with both CLIP (Radford et al., 2021) and SigLIP (Tschannen et al., 2025) architectures.
+Then, we evaluate the effectiveness of these retrievers with 17 widely used VLMs across the cVQA and cIC downstream tasks.
+More details are in Appendix I.
+3.1 Evaluation Metrics
+For multimodal retrieval, the performance is evaluated using standard retrieval metrics, including Mean Reciprocal Rank (MRR) (Shi et al., 2012), Precision@k (P@k) (Järvelin and Kekäläinen, 2017), and Normalized Discounted Cumulative Gain (nDCG@k) (Wang et al., 2013), where $k\in\{1,3,5\}$.
+Table 1: Kendall’s $\tau$ rank correlation (Kendall, 1938) between automatic metrics and human judgments for the cIC task.
+Statistically significant correlations ($p<0.05$) are marked with ✓.
+Our proposed metrics correlate stronger with human evaluation than the others.
+Rouge-L CIDER BERTScore CLIPScore RegionScore (ours)
+-0.172 ✗ -0.316 ✗ -0.011 ✗ 0.139 ✗ 0.442 ✓
+For cVQA, we use accuracy as the evaluation metric, as all tasks follow a multiple-choice format.
+Accuracy is computed per VLM and per retriever, representing the proportion of correctly answered questions.
+For cIC, we employ several metrics for evaluating the generated caption, including ROUGE-L (Lin, 2004), CIDEr (Vedantam et al., 2015), BERTScore (Zhang* et al., 2020), and CLIPScore (Hessel et al., 2021).
+To further evaluate the cultural relevance of the generated captions, we conducted a human study in which four researchers selected the most accurate captions from 14 VLMs.
+[^6] The Qwen3-VL models were omitted from the human evaluation for the cIC task because the used model variants were not released at that time.
+We find a significant mismatch between automatic metrics and human judgments of cultural appropriateness (Table 1).
+To bridge this gap, we further introduce RegionScore, a novel evaluation metric designed to quantify the extent to which captions reference specific geopolitical regions.
+See details in Appendix G and H.
+RegionScore.
+Let ${\mathbb{I}}=\{I_{1},I_{2},\dots,I_{n}\}$ denote the set of images.
+For each image $I_{i}$, let ${\bm{g}}^{(i)}$ denote its generated caption, and let ${\bm{c}}^{(i)}$ represent the country name associated with $I_{i}$.
+We define ${\mathbb{A}}_{i}$ as the set of adjectives or demonyms corresponding to ${\bm{c}}^{(i)}$.
+For convenience, we introduce the set of region-related terms associated with $I_{i}$: ${\mathbb{T}}_{I_{i}}=\{{\bm{c}}^{(i)}\}\cup{\mathbb{A}}_{i}$.
+The instance-level $R(\displaystyle{\bm{g}}^{(i)},I_{i})$ metric for sample $I_{i}$ and ${\bm{g}}^{(i)}$ is defined as:
+[…]
+The overall RegionScore for predicted captions are then computed as:
+$RegionScore_{\displaystyle{\mathbb{I}}}=\frac{1}{|\displaystyle{\mathbb{I}}|}\sum_{I_{i}\in\displaystyle{\mathbb{I}}}R(\displaystyle{\bm{g}}^{(i)},I_{i}),\quad RegionScore_{\displaystyle{\mathbb{I}}}\in[0,1]$ (2)
+RegionScore is designed to complement, rather than replace, existing general-purpose captioning metrics (e.g., CLIPScore), which primarily assess semantic alignment but do not explicitly capture geographic specificity.
+Therefore, RegionScore serves as a proxy for regional awareness by quantifying the proportion of captions that include explicit references of a country or its common demonyms, with higher values indicating stronger region awareness.
+In CCUB, $\text{RegionScore}_{\text{ground truth}}$ is 99%.
 3.2 Culture-aware Contrastive Learning
-Given an image $I_{i}$ associated with $T$ textual descriptions $\{D_{i1},D_{i2},\dots,D_{iT}\}$, each document $D_{it}$ is annotated with a binary label $y_{it}\in\{0,1\}$, where $y_{it}=1$ indicates cultural relevance and $y_{it}=0$ indicates irrelevance.
-For each image–text pair $(I_{i},D_{it})$, we employ a shared vision-language encoder—such as CLIP—to obtain modality-specific representations: $\mathbf{E}_{I_{i}}=\mathcal{E}_{V}(I_{i})$ for the visual input and $\mathbf{E}_{D_{it}}=\mathcal{E}_{L}(D_{it})$ for the textual input.
-We then compute the cosine similarity score $s_{it}$ between $\mathbf{E}_{I_{i}}$ and each corresponding $\mathbf{E}_{D_{it}}$, resulting in a similarity vector $\mathbf{S}_{i}=[s_{i1},s_{i2},\dots,s_{iT}]$.
+Given an image $I_{i}$ associated with a set of textual documents ${\mathbb{D}}=\{D_{i}^{1},D_{i}^{2},\dots,D_{i}^{n}\}$, each document $D_{i}^{j}$ is annotated with a binary label $y_{i}^{j}\in\{0,1\}$, where $y_{i}^{j}=1$ indicates cultural relevance and $y_{i}^{j}=0$ indicates irrelevance.
+For each image-document pair $(I_{i},D_{i}^{j})$, we employ a shared vision-language encoder, such as CLIP, to obtain modality-specific representations: $\mathbf{E}_{I_{i}}=\mathcal{E}_{V}(I_{i})$ for the visual input and $\mathbf{E}_{D_{i}^{j}}=\mathcal{E}_{L}(D_{i}^{j})$ for the textual input.
+We then compute the cosine similarity score $s_{i}^{t}$ between $\mathbf{E}_{I_{i}}$ and each corresponding $\mathbf{E}_{D_{i}^{j}}$, resulting in a similarity vector $\mathbf{S}_{i}=[s_{i}^{1},s_{i}^{2},\dots,s_{i}^{n}]$.
 Culture-awareness classification now amounts to:
-$\mathcal{L}_{\text{Culture Classify}}=-\frac{1}{B\cdot T}\sum_{i=1}^{B}\sum_{j=1}^{T}\left[y_{ij}\log\sigma(s_{ij})+(1-y_{ij})\log(1-\sigma(s_{ij}))\right],$ (1)
-where B is the number of images; $\sigma(\cdot)$ denotes the sigmoid function.
+$\mathcal{L}_{\text{Culture Classify}}=-\frac{1}{|{\mathbb{I}}|\cdot|{\mathbb{D}}|}\sum_{I_{i}\in{\mathbb{I}}}\sum_{D_{i}^{j}\in{\mathbb{D}}}\left[y_{i}^{j}\log\sigma(s_{i}^{j})+(1-y_{i}^{j})\log(1-\sigma(s_{i}^{j}))\right],$ (3)
+where $\sigma(\cdot)$ denotes the sigmoid function.
 To prioritize culturally relevant descriptions in the ranking, we apply a margin ranking loss between all pairs of descriptions with differing cultural relevance.
-For each image $I_{i}$, we compare all pairs $(D_{ij},D_{ik})$ such that $y_{ij}=1$ and $y_{ik}=0$, and encourage the model to assign a higher similarity score to the relevant description.
+For each image $I_{i}$, we compare all pairs $(D_{i}^{j},D_{i}^{k})$ such that $y_{i}^{j}=1$ and $y_{i}^{k}=0$, and encourage the model to assign a higher similarity score to the relevant description.
 The ranking loss is defined as:
-$\mathcal{L}_{\text{Rank}}=\frac{1}{B}\sum_{i=1}^{B}\sum_{\begin{subarray}{c}j,k=1\\ y_{ij}=1,y_{ik}=0\end{subarray}}^{T}\max\left(0,\delta-(s_{ij}-s_{ik})\right),$ (2)
+$\mathcal{L}_{\text{Rank}}=\frac{1}{|{\mathbb{I}}|}\sum_{i=1}^{|{\mathbb{I}}|}\sum_{\begin{subarray}{c}j,k=1,\\ y_{i}^{j}=1,y_{i}^{k}=0\end{subarray}}^{|{\mathbb{D}}|}\max\left(0,\delta-(s_{i}^{j}-s_{i}^{k})\right),$ (4)
 To mitigate the risk of overly similar positive text embeddings for the same image, we introduce a penalty that encourages intra-modal diversity among textual representations.
@@
 Specifically, the penalty is formulated using an exponential function to emphasize the dissimilarity between embeddings:
-$\mathcal{L}_{\text{Diversity}}\left(\mathbf{S}_{i}\right)=-\sum_{t=1}^{T}\log\left(\frac{\exp(s_{it})}{\sum_{j=1}^{T}\exp(s_{ij})}\right)$ (3)
+$\mathcal{L}_{\text{Diversity}}\left(\mathbf{S}_{i}\right)=-\sum_{t=1}^{|{\mathbb{D}}|}\log\left(\frac{\exp(s_{i}^{t})}{\sum_{j=1}^{|{\mathbb{D}}|}\exp(s_{i}^{j})}\right)$ (5)
 Then we can get the culture-aware contrastive loss:
@@
 Table 2: Performance with different retriever models.
-Fine-tuned contrastive models consistently outperform their frozen counterparts across tasks.
-Here, "CaSigLIP"=Culture-aware SigLIP2-SO/14@384px, and "CaCLIP"=Culture-aware CLIP-L/14@224px.
-Models in gray are frozen.
-Method MRR $\uparrow$ P@1 $\uparrow$ P@3 $\uparrow$ P@5 $\uparrow$ nDCG@1 $\uparrow$ nDCG@3 $\uparrow$ nDCG@5 $\uparrow$ SigLIP2-SO/14@384px 66.71 50.42 40.14 34.25 58.85 62.43 67.50 CLIP-L/14@224px 70.76 54.58 43.47 36.58 62.54 67.25 72.43 VisualBERT 59.66 42.50 35.97 32.75 51.29 55.49 62.29 VL-T5 55.53 35.42 34.03 31.42 45.21 53.18 59.62 LLaVA-OneVision-7B 54.15 36.20 31.83 28.69 45.85 50.82 56.74 CaSigLIP2 (ours) 69.35 54.17 43.47 36.50 61.98 66.75 71.98 CaCLIP (ours) 78.34 65.42 49.44 39.50 72.25 75.22 79.32
-4.2 Multimodal Retrieval Results
+Fine-tuned contrastive models consistently outperform their frozen counterparts (shown in gray).
+Ravenea-SigLIP and Ravenea-CLIP are SigLIP2-SO/14@384px and CLIP-L/14@224px models fine-tuned on Ravenea, respectively.
+Method MRR $\uparrow$ P@1 $\uparrow$ P@3 $\uparrow$ P@5 $\uparrow$ nDCG@1 $\uparrow$ nDCG@3 $\uparrow$ nDCG@5 $\uparrow$ SigLIP2-SO/14@384px 68.62 54.66 37.47 32.92 61.22 63.82 71.44 CLIP-L/14@224px 75.44 60.87 41.41 34.41 67.75 72.31 78.09 VisualBERT 61.33 46.71 33.14 29.42 55.76 59.42 65.92 VL-T5 58.33 39.86 32.58 29.35 47.74 57.29 62.12 LLaVA-OV-7B 58.85 37.48 30.68 28.21 48.59 51.80 60.34 Ravenea-SigLIP (ours) 70.95 57.14 40.99 33.29 63.86 68.31 73.92 Ravenea-CLIP (ours) 82.17 72.05 45.76 36.77 77.08 78.96 84.09
+4 Experimental Results
+4.1 Main Results
+Multimodal retrieval results.
 We perform a comprehensive evaluation of both frozen and fine-tuned retrievers, and present the results in Table 2.
 We find that fine-tuned models, particularly those based on contrastive learning, consistently outperform their frozen counterparts.
-For instance, CaCLIP achieves a substantial improvement in P@1, rising from 54.58% to 65.42%, and sets a new SOTA across all evaluation metrics.
-Although SigLIP2-SO/14@384px also benefits from fine-tuning, the performance gains are comparatively modest.
-In contrast, models such as LLaVA-OneVision-7B, VL-T5, and VisualBERT lag behind after fine-tuning, even underperforming relative to frozen baselines.
-This underperformance likely stems from the fact that models such as LLaVA-OneVision-7B and VisualBERT were originally pretrained for generative tasks with different objectives, whereas CLIP-L/14@224px and SigLIP2-SO/14@384px were explicitly trained for similarity-based alignment, providing them with a structural advantage in retrieval settings.
-5 Multimodal Retrieval-augmented Visual Culture Understanding
-We then evaluate the effectiveness of these retrievers with 14 SOTA VLMs, spanning a diverse set of architectures. We conduct experiments on two downstream tasks: cVQA and cIC, respectively.
-5.1 Experimental Setup
-Models.
-We benchmark open and closed-weight SOTA VLMs on Ravenea, leveraging various retrievers against non-RAG baselines, assessing retrieval effectiveness across models of different sizes.
-The open-weight models include LLaVA-OneVision-7B li2025llavaonevision , Pixtral-12B agrawal2024pixtral , Phi-4 Multimodal-Instruct abouelenin2025phi , Gemma3-4B-Instruct and 27B-Instruct team2025gemma , Qwen2.5-VL-Instruct (3B, 7B, 72B  [^3] Due to computational constraints, we use the quantized version Qwen2.5-VL-Instruct-72B-AWQ.bai2025qwen2 , InternVL3 (2B, 8B, 38B) zhu2025internvl3 , and Deepseek-VL2 variants (Tiny and Base) wu2024deepseek .
-For the closed models, we adopt GPT-4.1 achiam2023gpt (accessed on 2025/04/14)  [^4] Knowledge cutoff: June 1, 2024; https://platform.openai.com/docs/models/gpt-4.1.
-Evaluation metrics.
-Table 3: Kendall’s $\tau$ rank correlation kendall1938new between automatic metrics and human judgments for the CCUB task.
-Statistically significant correlations ($p<0.05$) are marked with ✓.
-Our proposed metrics correlate stronger with human evaluation than the others.
-Rouge-L lin-2004-rouge CIDER vedantam2015cider BERTScore Zhang*2020BERTScore: CLIPScore hessel-etal-2021-clipscore RegionScore (ours)
--0.172 ✗ -0.316 ✗ -0.011 ✗ 0.139 ✗ 0.442 ✓
-For the cVQA task, we use accuracy as the primary evaluation metric, which measures the proportion of correctly predicted answers.
-For the cIC task, we employ several evaluation metrics including ROUGE-L lin-2004-rouge , CIDEr vedantam2015cider , BERTScore Zhang*2020BERTScore: , and CLIPScore hessel-etal-2021-clipscore , to assess the alignment between generated and reference captions across lexical, syntactic, and embedding-based levels.
-To further evaluate the cultural relevance and human-perceived quality, we conduct a human evaluation study.
-We employed four researchers to select the most accurate caption from 14 VLMs (see details in Appendix I and L), and find a significant mismatch between automatic metric scores and human judgments of cultural appropriateness (see Table 3).
-To bridge this gap, we further introduce RegionScore, a novel evaluation metric designed to quantify cultural grounding (see details in Appendix H and Table 9).
-It measures how well captions identify the correct country names tied to cultural elements, adding geographic and cultural specificity in image captioning.
-5.2 Overall Performance
-We present the main results in Table 4.
+For instance, Ravenea-CLIP achieves a substantial improvement in P@1, rising from 60.87% to 72.05%, and sets a new SOTA across all evaluation metrics.
+Although SigLIP2-SO/14@384px (Tschannen et al., 2025) also benefits from fine-tuning, the performance gains are comparatively modest.
+In contrast, models such as LLaVA-OV-7B (Li et al., 2025), VL-T5 (Cho et al., 2021), and VisualBERT (Li et al., 2019; 2020) lag behind after fine-tuning, even underperforming relative to frozen baselines.
+This underperformance likely stems from the fact that models such as LLaVA-OV-7B and VisualBERT were originally pretrained for generative tasks with different objectives, whereas CLIP-L/14@224px (Radford et al., 2021) and SigLIP2-SO/14@384px were explicitly trained for similarity-based alignment, providing them with a structural advantage in retrieval settings.
+Table 3: cVQA and cIC Performance w/ and w/o RAG. Models in gray are frozen retrievers. The best results are in bold. VLMs augmented with finetuned retriever generally perform better.
+Open-source Proprietary Retriever Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-78B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OV-7B Qwen3-VL-2B Qwen3-VL-8B Qwen3-VL-32B GPT-4.1 $cVQA_{Accuracy\uparrow}$ W/O RAG 67.1 49.8 68.9 62.7 52.6 83.3 64.1 71.3 84.2 66.5 79.4 42.6 73.2 62.7 37.8 73.7 81.3 86.6 SigLIP2-SO/14@384px 71.0 48.8 76.1 63.2 67.5 82.3 71.3 73.7 86.1 67.5 77.5 64.6 75.1 44.5 67.9 75.6 83.3 82.8 CLIP-L/14@224px 71.9 50.2 77.5 63.6 66.0 82.8 73.7 78.5 84.2 69.4 76.6 67.0 76.6 44.5 68.9 76.6 83.3 83.3 Finetuned Models VisualBERT 67.6 48.8 71.8 61.7 64.1 78.5 66.5 67.5 82.3 63.6 75.1 61.2 68.4 37.3 60.8 73.2 83.7 85.2 VL-T5 66.1 46.6 72.7 58.4 65.1 78.5 65.1 67.9 80.4 58.4 75.6 58.9 62.7 35.4 60.3 75.6 78.9 83.7 LLaVA-OV-7B 67.9 50.1 66.0 60.3 67.0 81.8 66.5 66.0 84.2 65.1 75.1 60.8 68.9 37.3 59.8 77.0 83.7 84.7 Ravenea-SigLIP (ours) 72.6 48.8 76.1 68.4 67.9 81.3 73.2 75.6 85.6 70.3 79.4 69.4 75.6 47.8 68.9 77.0 84.2 85.6 Ravenea-CLIP (ours) 73.3 50.7 75.1 64.1 69.9 82.3 76.1 81.3 85.2 69.9 79.9 69.4 75.1 50.2 69.4 77.5 84.2 86.1 $cIC_{RegionScore\uparrow}$ W/O RAG 57.5 26.5 63.3 53.1 69.4 71.4 38.8 49.0 65.3 73.5 75.5 28.6 53.1 49.0 59.2 61.2 73.5 67.3 SigLIP2-SO/14@384px 54.8 30.6 59.2 51.0 63.3 65.3 44.9 53.1 57.1 65.3 69.4 46.9 53.1 42.9 49.0 59.2 59.2 61.2 CLIP-L/14@224px 65.1 38.8 69.4 59.2 71.4 69.4 65.3 65.3 71.4 73.5 73.5 61.2 69.4 51.0 63.3 67.3 71.4 65.3 Finetuned Models VisualBERT 57.4 36.7 57.1 53.1 61.2 69.4 44.9 59.2 61.2 63.3 69.4 42.9 57.1 44.9 55.1 65.3 69.4 65.3 VL-T5 57.0 34.7 61.2 55.1 65.3 73.5 46.9 53.1 61.2 61.2 67.3 44.9 59.2 38.8 53.1 63.3 67.3 63.3 LLaVA-OV-7B 56.8 38.8 63.3 51.0 63.3 69.4 44.9 57.1 59.2 65.3 71.4 40.8 55.1 46.9 49.0 61.2 67.3 61.2 Ravenea-SigLIP (ours) 60.6 30.6 73.5 57.1 67.3 71.4 44.9 59.2 69.4 67.3 69.4 57.1 61.2 40.8 53.1 65.3 71.4 71.4 Ravenea-CLIP (ours) 68.8 46.9 73.5 61.2 77.6 75.5 69.4 67.3 73.5 75.5 75.5 63.3 71.4 55.1 67.3 73.5 73.5 69.4
+Downstream tasks.
+We present downstream performance on both cVQA and cIC tasks in Table 3.
 The results demonstrate the efficacy of incorporating culture-aware retrieval augmentation.
-Employing fine-tuned retrievers yields substantial performance gains over both non-RAG and frozen retrievers baselines.
-Specifically, the CaCLIP achieves the highest average performance across both tasks, improves the accuracy for cVQA from 67.7% to 71.5%, and substantially improving the RegionScore for cIC from 40.2% to 58.1%.
-While CLIP-L/14@224px also offers improvements, fine-tuning consistently unlocks further potential.
-Furthermore, in the cVQA task, among all evaluated models, GPT-4.1 achieves the highest accuracy (86.8%) without RAG.
-Within the category of open-weight models, Qwen2.5-VL-72B leads with an accuracy of 81.0%.
-For lightweight models ($\leq$ 8B parameters), Qwen2.5-VL-7B achieves the best performance without RAG, reaching 67.7%.
-However, incorporating a CaCLIP significantly boosts performance, which enables InternVL3-8B to achieve 74.2% and outperform Qwen2.5-VL-7B by 0.6% with identical reranking.
-Notably, across multiple model families, augmenting the smallest variant with CaCLIP consistently elevates its performance to match or even exceed that of the next larger model tier.
-In the cIC task, with culture-aware contrastive learning, CaCLIP demonstrates substantial gains in identifying the culture in country-level of visual content, especially when built on top of VLMs with strong vision-language priors.
-It achieves the highest average RegionScore (58.1%) among the six reranking methods evaluated, with peak performance reaching 76.3% on the Gemma3-4B backbone.
-CaCLIP achieves leading the scores on 9 of 14 diverse suits of VLMs.
-This result underscores CaCLIP’s robustness and adaptability, particularly in culture-aware image captioning and retrieval tasks that demand fine-grained multimodal alignment.
-Table 4: cVQA and cIC Performance w/ and w/o RAG. Models in gray are frozen retrievers. Results are colored as Best. VLMs augmented with finetuned retriever generally perform better.
-Open Weights Closed Weights Retriever Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $cVQA_{Accuracy\uparrow}$ W/O RAG 67.7 57.1 65.5 64.2 67.7 81.0 58.4 66.5 71.9 63.6 78.7 55.5 66.1 64.8 86.8 SigLIP2-SO/14@384px 67.2 58.7 55.8 63.9 67.7 77.1 62.3 68.4 71.9 60.7 71.9 62.6 69.4 67.1 82.6 CLIP-L/14@224px 70.6 64.2 56.1 67.7 73.2 79.0 68.1 71.0 76.1 66.1 74.5 69.4 71.0 69.7 81.9 Finetuned Models VisualBERT 67.1 58.7 59.4 66.5 68.4 77.1 61.3 66.1 71.9 60.7 73.6 63.2 66.5 63.9 82.3 VL-T5 65.8 56.5 61.3 62.9 69.4 77.1 60.3 66.8 70.3 55.8 74.2 61.0 61.0 61.9 82.6 LLaVA-OneVision-7B 67.8 60.0 61.3 64.8 71.3 80.7 61.3 64.8 73.9 62.3 73.6 62.6 67.4 63.9 81.3 CaSigLIP2 (ours) 69.8 62.6 58.7 67.1 72.3 77.7 65.8 72.6 75.8 64.2 75.5 64.2 70.7 69.4 82.6 CaCLIP (ours) 71.5 65.8 60.0 69.4 73.6 81.0 68.7 74.2 75.2 66.8 75.2 66.8 71.0 70.0 83.2 $cIC_{RegionScore\uparrow}$ W/O RAG 40.2 31.3 36.3 36.3 48.8 47.5 27.5 37.5 22.5 67.5 62.5 10.0 35.0 18.8 47.5 SigLIP2-SO/14@384px 52.5 50.0 56.3 36.3 57.5 53.8 58.8 56.3 61.3 70.0 66.3 30.0 53.8 41.3 53.8 CLIP-L/14@224px 55.4 56.3 63.8 41.3 66.3 60.0 61.3 62.5 67.5 70.0 70.0 31.3 55.0 47.5 58.8 Finetuned Models VisualBERT 56.3 58.8 63.8 41.3 58.8 62.5 61.3 61.3 66.3 71.3 70.0 30.0 58.8 45.0 62.5 VL-T5 55.4 60.0 65.0 41.3 62.5 66.3 62.5 58.8 67.5 71.3 68.8 30.0 61.3 40.0 61.3 LLaVA-OneVision-7B 56.0 60.0 65.0 40.0 60.0 62.5 61.3 62.5 66.3 71.3 71.3 28.8 57.5 47.5 60.0 CaSigLIP2 (ours) 56.3 57.5 65.0 36.3 61.3 63.8 58.8 61.3 67.5 71.3 70.0 27.5 57.5 46.3 65.0 CaCLIP (ours) 58.1 60.0 67.5 42.5 65.0 65.0 62.5 60.0 70.0 76.3 75.0 27.5 62.5 52.5 55.0
-6 Analysis and Further Discussion
-Culture-aware retrieval augmentation substantially benefits VLMs across both cVQA and cIC tasks, compared to their no-RAG counterparts.
-In this section, we explore the margin of the improvement, cultural preference and the effectiveness of cultural annotation.
-6.1 Scaling Models Yields Diminishing or Negative Returns across Retrievers
-Figure 5: Performance improvements for smallest and largest models per family with multimodal retrievers.
-Scaling models yields marginal gains with various retrievers, even negative effects in both cVQA and cIC tasks.
-"ACC." denotes accuracy; "R.S." refers to the RegionScore; "$\Delta$" represents the change incorporated with RAG compared to the non-RAG baseline.
-In the cVQA task, within the same VLM family, performance differences between RAG and non-RAG approaches exhibit non-monotonic trends as model size scales.
-For all four model families, larger models show marginal or even negative returns from RAG integration.
-What’s more, sensitivity to RAG varies across model families.
-Notably, DeepSeek-VL2 demonstrates the most pronounced performance gap: the smallest model benefits from RAG with an average improvement of approximately +5%, whereas the largest model in the same family suffers a degradation of around -6% on average.
-In the cIC task, the effectiveness of RAG exhibits a consistent trend with respect to model scale within a given model family.
-Across all four model families evaluated, larger models tend to benefit less—or at most comparably—from the integration of RAG, suggesting diminishing returns at higher capacity.
-Among them, Gemma3 models show the smallest relative improvement, achieving approximately a +7% performance gain on average, whereas InternVL3 models yield the highest benefit, with performance gains reaching up to +30%.
-6.2 Differences across Countries
-We evaluate all models using CaCLIP across a range of countries for both tasks, as shown in Figure 6.
-In the cVQA setting, most VLMs exhibit substantially diminished performance on culture-specific questions regarding Nigeria and Indonesia, in contrast to their performance on questions under other national contexts.
-Interestingly, questions related to Spanish culture reveal high inter-model variance,
-Figure 6: Performance of 14 VLMs with CaCLIP across different countries. Despite the integration of CaCLIP, disparities in model performance persist across countries.
-with accuracy differentials reaching up to 35%, underscoring significant discrepancies in cultural representation across models.
-In the cIC task, VLMs consistently underperform on images and documents associated with Indian cultural contexts, while achieving the highest RegionScores on Korean culture-related inputs.
-Model performance on Indian culture is particularly volatile, indicating inconsistent cultural grounding across architectures.
+Our proposed fine-tuned retrievers, Ravenea-SigLIP and Ravenea-CLIP, consistently achieve the highest scores across both cVQA and cIC tasks, frequently outperforming both the non-RAG baselines and frozen retrievers (SigLIP2 / CLIP).
+Specifically, averaged over all VLMs, Ravenea-CLIP achieves the highest performance for both tasks, improving +6.2% in cVQA accuracy (67.1% $\rightarrow$ 73.3%), and +11.3% in cIC RegionScore (57.5% $\rightarrow$ 68.8%).
+While CLIP-L/14@224px also offers improvements, culture-aware contrastive learning consistently unlocks further potential.
+Furthermore, without retrieval, larger VLMs still achieve strong absolute performance on both tasks (e.g., InternVL3-78B at 84.2% on cVQA, and Gemma3-27B at 75.5% on cIC).
+Retrieval contents, however, disproportionately benefits small and mid size models.
+For lightweight models ($\leq$ 8B parameters), Ravenea-CLIP narrows the gap to the larger models.
+For example, Qwen3-VL-8B improves to 77.5% on cVQA, approaching the performance of the models an order of magnitude larger, highlighting the potential of culture-aware retrieval.
+4.2 Further Findings
+Culture-aware retrieval augmentation substantially benefits VLMs across both cVQA and cIC tasks, compared to their non-RAG counterparts.
+In this section, we explore the margin of improvement, cultural preference, the effectiveness of cultural annotation, and the impact of context length.
+Finding 1. While larger VLMs see limited gains from different retrievers, lightweight VLMs benefit substantially from culturally grounded retrieval contexts.
+Figure 4: Performance improvements per model family with Ravenea-CLIP.
+Scaling VLMs yields marginal gains on both cVQA and cIC tasks.
+$\Delta$" represents the change with Ravenea-CLIP relative to the non-RAG baseline.
+To evaluate the scaling behavior of VLMs incorporating with Ravenea-CLIP, we select four model families spanning from 2B to 78B parameters and quantify the performance improvement $\Delta$ relative to non-RAG baselines.
+As shown in Figure 4, the impact of Ravenea-CLIP does not scale monotonically with model size: smaller models exhibit the most substantial gains, while larger models (27B to 78B) exhibit only modest improvements.
+For instance, Qwen3-VL-2B achieves +31.6% absolute improvement on cVQA and +8.1% on cIC, whereas Qwen3-VL-32B shows only +2.9% better on cVQA and no gain on cIC.
+This trend indicates that larger models may already internalize much of the relevant knowledge, especially the Wikipedia content retrieved by Ravenea-CLIP, resulting in redundancy and a plateau in performance.
+Overall, these results suggest that Ravenea-CLIP serves as an effective plug-and-play enhancement for lightweight models on both cVQA and cIC tasks, providing diminishing yet consistent gains as model size increases.
+Finding 2. With Ravenea, VLMs show cultural disparities with and without RAG. Moreover, large inter-model variance highlights model-specific biases.
+Figure 5: Geographic variability in VLM performance. While Ravenea-CLIP yields gains in general, there’s still variance in both baseline and relative improvement across geographic regions.
+For geographic variability, we evaluate all models using Ravenea-CLIP across a range of countries for both tasks, as shown in Figure 5.
+For cVQA, most VLMs exhibit substantially diminished performance on culture-specific questions regarding Nigeria and Indonesia, in contrast to their performance on questions under other national contexts.
+Interestingly, questions related to Spanish culture reveal high inter-model variance, with accuracy differentials reaching up to 50%, underscoring significant discrepancies in cultural representation across models.
+For cIC, VLMs consistently underperform on images and documents associated with Mexico cultural contexts, while achieving the highest RegionScores on Indian culture-related inputs.
+Model performance on Mexico culture is particularly volatile, indicating inconsistent cultural grounding across architectures.
 By comparison, Korean and Chinese cultural inputs yield more stable performance across models, suggesting entrenched model-specific preferences in cultural alignment (see more results in Appendix J).
-6.3 Ablation Study on Annotation Questions
-Figure 7: Ablation for different annotation questions. Combining all three culture-relevant annotation questions yields the best performance.
+Finding 3.
+Generally, models benefit most when trained with the full set of culture-relevant annotations.
+Culturally grounded retrieval substantially outperforms random or irrelevant contexts, underscoring both the necessity of culture-aware contexts.
+Furthermore, more or longer retrieved context length yield diminishing gains.
+Figure 6: Ablations for annotations and retrieval content.
+Combining all three culture-relevant annotation questions consistently achieves the highest performance on both downstream tasks, validating the effectiveness of Ravenea setup.
+Annotation questions ablations.
 We further perform ablation studies across diverse combinations of annotation questions to assess their impact on downstream performance.
-Specifically, we evaluate 13 open-weight VLMs equipped with either CaSigLIP or CaCLIP, each trained on datasets constructed using varying subsets of culture-relevant annotations.
-From Figure 7, we can observe that leveraging all three questions (Q1 regarding country association; Q2 for topic alignment; Q3 for visual representation fidelity) yields the strongest performance on both cVQA and cIC tasks.
-For the cVQA task, we find Q1 provides the most significant benefit to CaSigLIP, whereas CaCLIP gains more from Q2.
-Among all pairwise combinations, the joint supervision from Q1 (country associations) and Q2 (topic alignment) proves slightly more effective than other pairs.
-In the cIC task, both CaSigLIP and CaCLIP achieve better performance improvements when trained with data derived from Q1, compared to other single-question sets.
-For pairwise combinations, CaCLIP benefits most from the Q1+Q3 combination, while CaSigLIP shows a clear preference for the Q2+Q3 setup.
+Specifically, we evaluate 16 open-weight VLMs equipped with either SigLIP2-SO/14@384px or CLIP-L/14@224px, each fine-tuned on datasets constructed using varying subsets of culture-relevant annotations.
+From Figure 6(a), we can observe that leveraging all three questions (Q1 regarding country association; Q2 for topic alignment; Q3 for explicit visual representation) yields the strongest performance on both cVQA and cIC tasks.
+For the cVQA task, we find Q1 provides the most significant benefit to Ravenea-SigLIP, whereas Ravenea-CLIP gains more from Q2.
+Among all pairwise combinations, the joint supervision from Q1 and Q2 proves slightly more effective than other pairs.
+In the cIC task, both Ravenea-SigLIP and Ravenea-CLIP achieve better performance improvements when trained with data derived from Q1, compared to other single-question sets.
+For pairwise combinations, Ravenea-CLIP benefits most from the Q1+Q3 combination, while Ravenea-SigLIP shows a clear preference for the Q2+Q3 setup.
+Context relevance ablations.
+Recent work (Roth et al., 2023; Ma et al., 2025) suggests that randomly augmented prompts may yield performance gains even in the absence of meaningful knowledge, raising concerns that improvements attributed to retrieval could in fact stem from other factors, such as superficial prompt lengthening and so on.
+To rigorously assess the effectiveness of culturally grounded retrieval, we design two control conditions: random character sequences, culturally unrelated documents.
+For random character sequences, we randomly generate synthetic strings in which each string consists of 100 to 300 substrings, and each substring contains 4 to 8 randomly sampled alphabetic characters, ensuring comparable token budgets without introducing interpretable content.
+For culturally unrelated documents, we randomly sample hundreds of Wikipedia articles from countries not represented in Ravenea, thereby maintaining language structure and topical coherence while explicitly breaking cultural relevance.
+Together, these conditions isolate whether gains arise from the related cultural context or merely from the presence of additional text.
+As shown in Figure 6(b), models augmented with Ravenea-CLIP consistently outperform other configurations, highlighting both the effectiveness of cultural retrieval and the robustness of the Ravenea evaluation protocol.
+More results are provided in Appendix J Table 11.
+Figure 7: Context length ablations for cVQA task. With Ravenea-CLIP, adding longer or multiple documents yields diminishing returns, highlighting the importance of the context length.
+Context length ablations.
+We also extend our study with a controlled analysis of retrieval length, comparing three additional context granularities: paragraph-level, full top-1 document, and full top-3 documents.
+As shown in Figure 7, providing a compact, high-salience context (Top-1 256 tokens) consistently achieves the best or tied-best performance across most model scales.
+In contrast, expanding the context to full documents or multiple documents often leads to marginal gains or degradation.
+This trend is particularly evident for lightweight models, which appear more sensitive to the context length.
+Even for larger models, additional context yields only marginal improvements despite substantially increasing input length.
+These results indicate that simply retrieving more content does not necessarily translate into better performance, especially for the lightweight models.
+5 Related Work
+Retrieval augmentation for cultural understanding.
+Retrieval augmentation has demonstrated significant efficacy for culture-related NLP tasks (Conia et al., 2024; Hu et al., 2024; Fung et al., 2023).
+Some works have explored retrieval-augmented approaches to cultural question answering.
+One line of work implemented retrieval using the World Values Survey to enable in-context learning for culturally contextualized questions (Seo et al., 2025; Yadav et al., 2025b), while others enhanced cultural question answering by retrieving information from web searches and knowledge bases (Lertvittayakumjorn et al., 2025).
+In cross-cultural applications, retrieval has also been shown effective for adapting recipes across cultural contexts (Hu et al., 2024).
+For multimodal understanding, however, the direction of culture-relevant retrieval remains underexplored.
+Vision-language culture datasets.
+Recent studies examine cultural understanding in VLMs through multicultural VQA (Liu et al., 2021; Nayak et al., 2024; Liu et al., 2025).
+Some researchers focus on culturally grounded visual categories such as cuisine, where models must distinguish dishes that are visually similar yet culturally distinct, highlighting fine-grained, geographically situated knowledge gaps (Winata et al., 2025; Li et al., 2024).
+Socioeconomic signals have also been explored, for example by studying how models infer or misinfer income level and living standards from visual cues (Nwatu et al., 2025).
+Others benchmark cultural entity recognition using Wikipedia-based prompts (Nikandrou et al., 2025), curate culture-specific image sets for value-based tasks (Yadav et al., 2025b) or concept-based image retrieval without explicitly considering cultural relevance (Bhatia et al., 2024).
+While prior efforts highlight the limitations of current VLM, they mainly assess the performance on retrieval and downstream tasks respectively, which are not suitable for probing how RAG shapes visual-cultural understanding.
+Our work specifically aims to fill this research gap.
 6 Conclusion
-We introduce Ravenea, a novel benchmark dataset designed to comprehensively evaluate the cultural sensitivity of diverse multimodal retrievers across 14 SOTA VLMs.
-The benchmark consists of culturally contextualized queries and image captions from eight countries, curated and paired with Wikipedia passages ranked by human annotators based on their cultural relevance.
-Our findings highlight the potential of RAG for visual cultural understanding, particularly when lightweight VLMs are enhanced with culturally-aware multimodal retrievers such as CaCLIP, which consistently outperform their non-augmented counterparts – with the caveat that the largest models in each VLM family often exhibit diminishing returns when integrated with RAG.
-Notably, the inclusion of culture-sensitive questions during data annotation significantly improves effectiveness of multimodal retrievers and enhances performance in downstream tasks.
-8 Acknowledgment
-We would like to thank all the annotators for their work.
-Special thanks to Nico Lang, Peter Ebert Christensen, Zhaochong An, Stella Frank, Srishti Yadav, for providing helpful research advice.
-Serge Belongie and Jiaang Li are supported by the Pioneer Centre for AI, DNRF grant number P1.
-Ivan Vulić is supported by a personal Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137).
-Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145).
-Appendix A Limitations
-While Ravenea establishes a solid foundation for advancing the study of visual culture understanding with retrieval augmentation, it has three limitations that warrant future attention.
-First, due to budgetary constraints, the dataset’s cultural scope is currently limited to eight countries and eleven categories.
-Although this selection introduces meaningful diversity, it does not comprehensively represent the global spectrum of cultural perspectives, particularly those of underrepresented or marginalized communities.
-Second, our use of Wikipedia as the primary external knowledge source introduces inherent biases and may lack the depth, plurality, and contextual richness necessary for nuanced cultural interpretation.
-Finally, due to resource limitations, we were unable to include certain proprietary VLMs that require paid APIs, such as Gemini 2.5 Pro [52] and Claude Opus 3.7 [53].
-We hypothesize that their performance would be comparable to GPT-4.1, which was included in our evaluation, but this remains an open empirical question.
-Appendix B Future Directions
-Our work opens several avenues for advancing visual culture understanding in multimodal models.
-First, expanding Ravenea to include more countries, cultural categories, and diverse knowledge sources, beyond Wikipedia, would improve coverage and reduce institutional bias.
-Second, future benchmarks could include richer tasks beyond cVQA and cIC, such as culture-grounded object recognition, historical retrieval, and symbolic interpretation, to better capture cultural semantics.
-Third, our results suggest a need for culturally-aware evaluation metrics, particularly for text generation.
-The limited effectiveness of retrieval augmentation in larger models also warrants further study, especially regarding how cultural knowledge is integrated and utilized.
-Together, these directions aim to support the development of more culturally-sensitive and globally robust vision-language models.
-Appendix C Ethics Statement
+We presented Ravenea, a new benchmark for multimodal retrieval-augmented visual culture understanding that enables the evaluation of how external cultural knowledge improves VLMs.
+By pairing culture-centric visual question answering and culture-informed captioning with 11,396 human-ranked Wikipedia documents across eight countries and eleven categories, Ravenea provides the first large-scale testbed for studying multimodal retrieval in culturally grounded settings.
+Our experiments across seven multimodal retrievers and seventeen VLMs show that incorporating cultural grounding consistently improves both retrieval quality and downstream task performance.
+Notably, lightweight VLMs benefit significantly from culture-aware retrieval, highlighting multimodal retrieval augmentation as an efficient way to enhance culturally situated understanding and reasoning.
+Overall, Ravenea offers a standardized framework for diagnosing limitations and advancing research in multimodal retrieval, grounding, and generation for VLMs in culture domain.
+We hope this benchmark will drive future work toward more culture-sensitive and globally inclusive multimodal intelligence.
+Ethics Statement
 This work focuses on improving the cultural awareness of VLMs through retrieval-augmented methods.
@@
 Consequently, the benchmark cannot capture the full richness of any cultural context.
+Unavoidably, when models fail, errors may go beyond reduced accuracy, leading to misattributed traditions or distorted cultural practices.
+In heritage contexts, such failures may undermine trust in digital tools or inadvertently disrespect the communities represented.
+Reliance on broad but uneven sources like Wikipedia can further compound these risks by amplifying existing biases and privileging cultures with richer digital footprints, echoing broader challenges in AI for cultural heritage (Münster et al., 2024).
+Such distortions could affect heritage management by marginalizing underrepresented voices, eroding local agency, or encouraging overreliance on automated outputs.
+To mitigate these risks, we emphasize safeguards such as human-in-the-loop validation with cultural experts, transparent documentation of dataset limitations, and fairness-oriented evaluation protocols.
+Looking forward, retrieval grounded in domain-specific knowledge bases curated by GLAM (Galleries, Libraries, Archives, and Museums) institutions offers a path toward higher-quality cultural data.
+We also highlight the importance of systematically incorporating expert human evaluation to validate both cultural relevance and interpretive accuracy.
 Finally, this work does not involve any personally identifiable information (PII), biometric data, or sensitive attributes.
 All human annotators were compensated fairly, and data annotation adhered to ethical guidelines for responsible research.
-Appendix D Experimental Details for Multi-modal Retrievers
-Dataset.
-We integrate responses from three annotation questions per data point into a continuous scale ranging from $-3$ to $3$, where higher values indicate stronger cultural relevance.
-To ensure fair evaluation across regions, we adopt the train-validation-test split strategy from class-imbalance loss [54], yielding an approximate 85-5-10 data split.
-This setup guarantees that validation and test sets contain a balanced set of images per country (Figure 8), thereby mitigating evaluation bias from skewed geographic distributions.
-During training, we employ random cropping for images as a sample-level augmentation to normalize the distribution of training instances across countries, further mitigating region-specific sampling biases.
-Figure 8: Data distributions across eight countries. Restrict the dataset to the images paired with at least one document annotated by human raters as culturally relevant.
-Table 5: Hyperparameters for finetuning on five models.
-Hyperparameters VisualBERT VL-T5 LLaVA-OneVision-7B CLIP-L/14@224px SigLIP2-SO/14@384px
-batch size 32 128 4 64 64
-lr 2e-4 1e-4 1e-4 1e-5 1e-5
-lr warmup ratio - 0.1 - - -
-weight decay - - - - -
-Max Epoch 100 20 10 50 50
-Patience 10 10 3 5 5
-early stopping Yes
-optimizer AdamW
-Using LORA - - Yes - -
-Hyperparameters.
-In this work, we adopt different sets of hyperparameters as VisualBERT, VL-T5, LLaVA-OneVision-7B, CLIP-L/14@224px, SigLIP2-SO/14@384px.
-For VisualBERT, VL-T5, and LLaVA-OneVision-7B we follow the setting in [44, 45, 46].
-We show the training hyperparameters in rerankering experiments for all models in Table 5.
-All experiments are conducted using a maximum of 2 Nvidia H100 GPUs.
-Appendix E Data Statistics
-Table 6: Statistics of images in each country.
-India Mexico Indonesia China Nigeria Korea Spain Russia
-408 349 309 303 223 151 142 77
-Table 7: Statistics of images in each category.
-Architecture Cuisine History Art Daily Life Companies Sports & Recreation Transportation Religion Nature Tools
-403 402 278 275 185 80 73 68 52 31 21
+Acknowledgment
+We would like to thank all the annotators for their work.
+Special thanks to Nico Lang, Peter Ebert Christensen, Zhaochong An, Stella Frank, Srishti Yadav, for providing helpful research advice.
+This project is mainly supported by the National Research Foundation Singapore under the AI Singapore Programme (AISG Award No: AISG3-RPGV-2025-016).
+Jiaang Li and Serge Belongie are supported by the Pioneer Centre for AI, DNRF grant number P1.
+Ivan Vulić is supported by a personal Royal Society University Research Fellowship ‘Inclusive and Sustainable Language Technology for a Truly Multilingual World’ (no 221137).
+Wenyan Li is supported by the Lundbeck Foundation (BrainDrugs grant: R279-2018-1145).
+Yang Deng is support by the Lee Kong Chian Fellowship awarded by Singapore Management University.
+Reproducibility Statement
+Section 2 and Appendix E, F and L details the data collection and processing steps. Experimental setups are provided in Appendix I. Code and dataset are available here.
+Appendix A The usage of Large Language Models (LLMs)
+We use LLMs as writing aids, limited to polishing grammar, improving clarity, and refining readability.
+Moreover, during data curation, we used LLMs to generate culture-informed captions solely as auxiliary references for annotators; these LLM outputs were not directly included in the final dataset.
+All technical content, results, and conclusions are original and authored entirely by the listed contributors.
+The responsibility for the manuscript’s scientific validity and integrity rests fully with the authors.
+Appendix B Limitations
+While Ravenea establishes a solid foundation for advancing the study of visual culture understanding with retrieval augmentation, it has three limitations that warrant future attention.
+First, due to budgetary constraints, the dataset’s cultural scope is currently limited to eight countries and eleven categories.
+Although this selection introduces meaningful diversity, it does not comprehensively represent the global spectrum of cultural perspectives, particularly those of underrepresented or marginalized communities.
+Second, our use of Wikipedia as the primary external knowledge source introduces inherent biases and may lack the depth, plurality, and contextual richness necessary for nuanced cultural interpretation.
+Finally, due to resource limitations, we were unable to include certain proprietary VLMs that require paid APIs, such as Gemini 2.5 Pro (Team et al., 2023) and Claude Opus 3.7 (Anthropic, 2024).
+We hypothesize that their performance would be comparable to GPT-4.1 (Achiam et al., 2023), which was included in our evaluation, but this remains an open empirical question.
+Appendix C Future Directions
+Our work opens several avenues for advancing visual culture understanding in multimodal models.
+In constructing Ravenea, we algorithmically generated a raw dataset spanning 30 countries, comprising over 3,900 images and 78,000 documents (Table 4).
+Due to budget constraints, annotations were limited to data from eight countries.
+Therefore, extending Ravenea to include more countries, cultural categories, and diverse knowledge sources, beyond Wikipedia, would improve coverage and reduce institutional bias.
+Second, future benchmarks could include richer tasks beyond cVQA and cIC, such as culture-grounded object recognition, historical retrieval, and symbolic interpretation, to better capture cultural semantics.
+Third, our results suggest a need for culturally-aware evaluation metrics, particularly for text generation.
+The limited effectiveness of retrieval augmentation in larger models also warrants further study, especially regarding how cultural knowledge is integrated and utilized.
+Exploring retrieval pipelines grounded in domain-specific knowledge bases curated by GLAM (Galleries, Libraries, Archives, and Museums) institutions could provide higher-quality cultural data and more reliable grounding.
+Finally, systematic human evaluation by cultural experts remains indispensable for validating both relevance and interpretive accuracy.
+Together, these directions aim to support the development of more culturally-sensitive and globally robust vision-language models.
+Appendix D Data Statistics
+Table 4: Statistics of the Ravenea dataset.
+The dataset is constructed by curating existing sources and augmenting them with wiki-derived documents to broaden cultural knowledge coverage and enhance content diversity.
+Dataset Countries Images Documents Pairs Questions Captions
+CVQAselected 7 1,213 8,194 12,130 2,331 -
+CCUBselected 5 655 4,354 6,550 - 655
+Ravenea 8 1,868 11,396 18,680 2,331 655
+BM25 filtered 30 3,912 78,240 - - -
+Table 5: Statistics of images in each country in Ravenea.
+India Mexico Indonesia Nigeria China Korea Spain Russia
+408 347 309 223 214 148 142 77
+Table 6: Statistics of images in each category in Ravenea.
+Architecture Cuisine Daily Life History Art Companies Sports & Recreation Transportation Religion Nature Tools
+403 402 347 278 86 80 73 68 52 31 21
 We then apply cryptographic hashing (SHA-256) to identify and remove duplicate images, resulting in a cleaner and more distinct cultural image set.
 Consequently, Ravenea comprises images collected from eight countries across four continents, spanning eleven distinct categories.
-The distribution of images by country and by category is detailed in Tables 6 and 7, respectively.
-Appendix F Evaluation of VLMs
-For closed-weight model, GPT-4.1, we directly call the corresponding API.
-For open-source models, we use vllm [55].
-During the evaluation, to ensure the stability of the results, we set the temperature parameter to 0.0 and the maximum output length to 256.
-All open-weight models are listed in Table 8.
-To avoid the impact of the length the retrieved content, we use the first 256 words in the top-1 Wikipedia document.
-Table 8: Model details: Hugging Face model names.
-Model Hugging Face Model Name
-LLaVA-OneVision-7B [19] llava-hf/llava-onevision-qwen2-7b-ov-hf
-Phi-4-Multimodal [21] microsoft/Phi-4-multimodal-instruct
-Pixtral [20] mistral-community/pixtral-12b
-Qwen2.5VL family [23]
-Qwen/Qwen2.5-VL-3B-Instruct
-Qwen/Qwen2.5-VL-7B-Instruct
-Qwen/Qwen2.5-VL-72B-Instruct-AWQ
-DeepSeek-VL2 family [25]
-deepseek-ai/deepseek-vl2
-deepseek-ai/deepseek-vl2-tiny
-InternVL3 family [24]
-OpenGVLab/InternVL3-2B
-OpenGVLab/InternVL3-8B
-OpenGVLab/InternVL3-38B
-Gemma3 family [22]
-google/gemma-3-4b-it
-google/gemma-3-27b-it
-Appendix G Annotation Details
+The distribution of images by country and by category is detailed in Tables 5 and 6, respectively.
+During constructing Ravenea, we have constructed a comprehensive, algorithmically generated raw dataset covering 30 countries, comprising over 3,900 images and 78,000 documents (see Table 4).
+Owing to budgetary limitations, for Ravenea, we have only annotated data from eight of these countries.
+Overall, we believe that both the raw algorithmically generated data and our human-labeled Ravenea could be a valuable source for improving the cultural understanding in VLM tasks.
+Appendix E Annotation Details
 Based on the initial BM25 retrieval results, we refine the cultural relevance label of retrieved documents via human annotation.
 We found that directly asking annotators to rate overall cultural relevance on a continuous scale (e.g., 0–10) led to unreliable and inconsistent labels.
-This difficulty arises from several factors: (1) the semantic meaning of intermediate scores is ambiguous, (2) annotators tended to overemphasize a few salient visual elements [56], and (3) small numerical differences (e.g., between 5 and 6) often fail to reflect meaningful distinctions, especially given the cognitive load of processing lengthy Wikipedia documents, resulting in intra-annotator variance even on repeated examples.
+This difficulty arises from several factors: (1) the semantic meaning of intermediate scores is ambiguous, (2) annotators tended to overemphasize a few salient visual elements (Chen et al., 2015), and (3) small numerical differences (e.g., between 5 and 6) often fail to reflect meaningful distinctions, especially given the cognitive load of processing lengthy Wikipedia documents, resulting in intra-annotator variance even on repeated examples.
 Instead, given an image–caption–document triplet, we decomposed cultural relevance into three interpretable and independently verifiable dimensions: country association, topic alignment, and explicit visual representation.
+Figure 8: Statistics on the Yes / No response distributions for the annotation questions.
 Prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines.
@@
 They are asked to assess whether each article provides meaningful background or contextual information that is directly relevant to the cultural elements described in the caption or depicted in the image.
-Appendix H A New Metric: RegionScore
-To evaluate the extent to which captions reference specific geopolitical regions, we introduce a RegionScore.
-This metric measures whether a caption contains explicit references to a country or its common demonyms.
-$N$ be the total number of samples.
-$\text{PRED}_{i}$ denote predicted captions for the $i$-th sample.
-$C_{i}$ be the country name associated with the $i$-th sample.
-$\text{Adj}(C_{i})$ denote the set of adjectives or demonyms associated with $C_{i}$.
-$\mathcal{T}_{i}=\{C_{i}\}\cup\text{Adj}(C_{i})$ denote the set of region-related terms for the $i$-th sample.
-We define binary indicators for each sample:
-$\delta^{\text{pred}}_{i}=\begin{cases}1&\text{if any }t\in\mathcal{T}_{i}\text{ appears in }\text{PRED}_{i}\\ 0&\text{otherwise}\end{cases}$ (5)
-The RegionScore for predicted captions are then computed as:
-$\text{RegionScore}_{\text{PRED}}=\frac{1}{N}\sum_{i=1}^{N}\delta^{\text{pred}}_{i}$ (6)
-These scores reflect the proportion of captions that include explicit regional identifiers.
-A higher score indicates stronger region-awareness in the captioning.
-For the ground truth in CCUB dataset, the $\text{RegionScore}_{\text{GT}}=99\%$.
-Appendix I Details of Human Evaluation in the cIC Task
-We randomly sample 10 images and generate captions using 14 vision-language models (VLMs) under three configurations: CaCLIP, CLIP, and a no-retrieval baseline, yielding 420 captions in total.
+The statistics of the annotations are presented in Figure 8.
+Table 7: Statistics of missing documents, along with the positive and negative ratios for each country.
+Country Positive & Negative Ratio(%) Missing Ratio(%)
+India 54.42 / 45.58 9.56
+Russia 54.42 / 45.58 0.0
+Spain 48.17 / 51.83 0.0
+Nigeria 34.47 / 65.53 0.45
+Mexico 38.37 / 61.63 25.55
+Indonesia 41.26 / 58.74 0.32
+Korea 69.59 / 30.41 19.59
+China 56.12 / 43.88 10.75
+As shown in Table 7, regions such as Mexico (25.55%) and Korea (19.59%) exhibit notably high missing ratios, underscoring the indispensable role of human expertise in identifying culturally specific documents that purely automated retrieval methods fail to capture.
+This missing ratio serves as an informative diagnostic metric: higher values directly reflect the effectiveness of our human-in-the-loop pipeline in enriching the dataset with highly local and otherwise overlooked content, ultimately attesting to the superior cultural coverage and quality of Ravenea.
+Furthermore, Table 7 reports the distribution of positive versus negative labels for document relevance across regions, offering additional insight into cultural representation variance.
+Appendix F Quality Control
+To evaluate the quality of the GPT-4o-generated captions, we randomly sampled 50 captions from the subset of Chinese images, and employ two local checkers to check the quality in Table 8.
+Each annotator assessed the accuracy of the captions, and we report their individual accuracies along with the IAA to measure consistency.
+Beyond objective correctness, we also asked annotators to provide subjective ratings (5-point Likert scale) of caption quality along two dimensions: naturalness (fluency and readability) and completeness (coverage of visual content).
+The results further show the GPT-4o caption is suitable for the next step to retrieve the Wikipedia documents.
+Table 8: Generated caption quality check. IAA denotes inter-annotator agreement.
+Acc. (Annotator 1) Acc. (Annotator 2) IAA Naturalness (1–5) Completeness (1–5)
+92% 94% 0.85 4.98 4.66
+To ensure the quality and consistency of our annotations, we implement several quality control methods.
+First of all, prior to the annotation process, all annotators are required to carefully review a detailed instruction file outlining the relevance criteria and annotation guidelines.
+To ensure proper understanding of the guidelines, annotators are required to complete a mock annotation test and correctly answer all questions before proceeding with the actual annotation tasks (see Figure 13).
+We also perform an additional quality check on a subset of the dataset.
+Specifically, for each selected countries, we employ an additional local quality checker who is tasked with manually reviewing the annotations to verify their accuracy and adherence to the guidelines.
+The quality checker reviews a random sample of annotated items, focusing on both the relevance labels and the justification behind any edge cases, such as borderline relevance or use of the “Cannot be determined” label.
+If inconsistencies or deviations from the annotation guidelines are identified, the affected samples are flagged for re-annotation.
+The overall acceptance rate from the meta quality checkers is 98.2%.
+The inter-annotator agreement (IAA) Cohen’s Kappa ($\kappa$) between the meta checker and annotator on the sampled annotations is 0.83.
+Appendix G Correlation between Automatic Metrics and Human Judgments
+To assess the correlation between human preferences and automatic evaluation metrics, we compute Kendall’s $\tau$ rank correlation.
+Human annotations are segmented according to the output chunks produced by corresponding 14 VLMs (no Qwen3VL family) with each retriever variant (Ravenea-CLIP-based, CLIP-based, and non-RAG).
+Within each segment, we calculate the selection winning ratio for each retrieval method, yielding a human preference vector formed by concatenating these ratios across all evaluation instances.
+For the automatic evaluation, we extract BERTScore, CIDEr, ROUGE-L, and CLIPScore for each corresponding retrieval variant and VLM.
+These scores are similarly concatenated into a metric-based vector.
+Finally, we compute Kendall’s $\tau$ between the human and metric vectors to quantify the consistency between automatic rankings and human judgments.
+Appendix H Details of Human Evaluation in the cIC Task
+We randomly sample 10 images and generate captions using 14 VLMs under three configurations: Ravenea-CLIP, CLIP, and a no-retrieval baseline, yielding 420 captions in total.
+[^7] The variants of the Qwen3-VL models we used were not released at the time of this evaluation.
 Four expert annotators participated in the evaluation.
@@
 Inter-annotator agreement (IAA) between this annotator and the original four, measured using Cohen’s $\kappa$, is 0.595.
-Appendix J Additional Results
-Figure 9: Performance of the 14 VLMs equipped with CaCLIP in cVQA task.
-As illustrated in Figures 9 and 10, the models exhibit diverse cultural preferences.
-Notably, most models achieve relatively stronger performance on Chinese and Indian cultural contexts in the cVQA task, and on Chinese and Korean contexts in the cIC task.
-Figure 10: Performance of the 14 VLMs equipped with CaCLIP in cIC task.
-We report the CIDEr, ROUGE-L, BERTScore, and CLIPScore metrics for 14 VLMs on the cIC task in Table 9.
+Appendix I Implementation Details
+I.1 Dataset.
+Figure 9: Data distributions across eight countries. Restrict the dataset to the images paired with at least one document annotated by human raters as culturally relevant.
+We integrate responses from three annotation questions per data point into a continuous scale ranging from 3-3 to $3$, where higher values indicate stronger cultural relevance.
+To ensure fair evaluation across regions, we adopt the train-validation-test split strategy from class-imbalance loss (Cui et al., 2019), yielding an approximate 85-5-10 data split.
+We also performed strict de-duplication on the training and validation sets, guaranteeing that no identical documents appear across the training, validation, testing splits.
+This setup guarantees that validation and test sets contain a balanced set of images per country in per task (Figure 9), thereby mitigating evaluation bias from skewed geographic distributions.
+During training, we employ random cropping for images as a sample-level augmentation to normalize the distribution of training instances across countries, further mitigating region-specific sampling biases.
+I.2 Multimodal Retrievers.
+We finetune CLIP-L/14@224px (Radford et al., 2021), SigLIP2-SO/14@384px (Tschannen et al., 2025).
+We also finetune a VisualBERT-based (Li et al., 2019; 2020) reranker following standard BERT-style setups (Nogueira and Cho, 2019), and adapt two multimodal generative models, such as VL-T5 (Cho et al., 2021) and LLaVA-OneVision-7B (Li et al., 2025), for end-to-end document retrieval (Chen et al., 2022; Yuan et al., 2024; Sun et al., 2023).
+We integrate responses from three annotation questions per data point into a continuous scale ranging from 3-3 to $3$, where higher values indicate stronger cultural relevance.
+Table 9: Hyperparameters for finetuning on five models.
+Hyperparameters VisualBERT VL-T5 LLaVA-OV-7B CLIP-L/14@224px SigLIP2-SO/14@384px
+batch size 32 128 4 64 64
+lr 2e-4 1e-4 1e-4 1e-5 1e-5
+lr warmup ratio - 0.1 - 0.03 0.03
+Max Epoch 100 20 10 2 2
+Patience 10 10 3 - -
+early stopping Yes Yes Yes No No
+optimizer AdamW
+Using LORA No No Yes No No
+Hyperparameters.
+In this work, we adopt different sets of hyperparameters as VisualBERT, VL-T5, LLaVA-OV-7B, CLIP-L/14@224px, SigLIP2-SO/14@384px.
+For VisualBERT, VL-T5, and LLaVA-OV-7B we follow the setting in (Chen et al., 2022; Yuan et al., 2024; Sun et al., 2023).
+We show the training hyperparameters in rerankering experiments for all models in Table 9.
+All experiments are conducted using a maximum of 2 Nvidia H100 GPUs.
+I.3 Vision-language models
+Table 10: Model details: Hugging Face model names.
+Model Hugging Face Model Name
+LLaVA-OneVision-7B (Li et al., 2025) llava-hf/llava-onevision-qwen2-7b-ov-hf
+Phi-4-Multimodal (Abouelenin et al., 2025) microsoft/Phi-4-multimodal-instruct
+Pixtral (Agrawal et al., 2024) mistral-community/pixtral-12b
+Qwen2.5VL family (Bai et al., 2025b)
+Qwen/Qwen2.5-VL-3B-Instruct
+Qwen/Qwen2.5-VL-7B-Instruct
+Qwen/Qwen2.5-VL-72B-Instruct-AWQ
+Qwen3VL family (Bai et al., 2025a)
+Qwen/Qwen3-VL-2B-Instruct
+Qwen/Qwen3-VL-8B-Instruct
+Qwen/Qwen3-VL-32B-Instruct-FP8
+DeepSeek-VL2 family (Wu et al., 2024)
+deepseek-ai/deepseek-vl2
+deepseek-ai/deepseek-vl2-tiny
+InternVL3 family (Zhu et al., 2025)
+OpenGVLab/InternVL3-2B
+OpenGVLab/InternVL3-8B
+OpenGVLab/InternVL3-78B-AWQ
+Gemma3 family (Team et al., 2025)
+google/gemma-3-4b-it
+google/gemma-3-27b-it
+We benchmark open and closed-weight widely used VLMs on Ravenea, leveraging various retrievers against non-RAG baselines, assessing retrieval effectiveness across models of different sizes.
+The open-sourced models include LLaVA-OneVision-7B (Li et al., 2025), Pixtral-12B (Agrawal et al., 2024), Phi-4 Multimodal-Instruct (Abouelenin et al., 2025), Gemma3-4B-Instruct and 27B-Instruct (Team et al., 2025), Qwen2.5-VL-Instruct (3B, 7B, 72B) (Bai et al., 2025b), Qwen3-VL-Instruct(3B, 7B, 32B), InternVL3 (2B, 8B, 78B) (Zhu et al., 2025), and Deepseek-VL2 variants (Tiny and Base) (Wu et al., 2024).
+For the closed models, we adopt GPT-4.1 (Achiam et al., 2023) (accessed on 2025/04/14).
+[^8] Knowledge cutoff: June 1, 2024; https://platform.openai.com/docs/models/gpt-4.1.
+For proprietary model, GPT-4.1, we directly call the corresponding API.
+For open-sourced models, we use vllm (Kwon et al., 2023).
+During the evaluation, to ensure the stability of the results, we set the temperature parameter to 0.0 and the maximum output length to 2048.
+All open-weight models are listed in Table 10.
+To avoid the impact of the length the retrieved content, we use the first 256 words in the top-1 Wikipedia document.
+Appendix J More Results
+Figure 10: Average improvement across 16 open-source VLMs for different countries with 2 retrievers on both cIC RegionScore and cVQA accuracy.
+(a) Performance of the 16 VLMs equipped with Ravenea-CLIP in cVQA task.
+(b) Performance of the 16 VLMs equipped with Ravenea-CLIP in cIC task.
+Figure 11: Performance of the 16 VLMs equipped with Ravenea-CLIP across tasks and countries.
+As illustrated in Figures 11(a) and 11(b), the models exhibit diverse cultural preferences.
+Notably, most models achieve relatively stronger performance on Indian cultural contexts in both cIC and cVQA tasks.
+As shown in Figures 10, VLMs demonstrate varying degrees of performance shifts across cultural contexts in both the cIC and cVQA tasks.
+On cIC, Ravenea-CLIP yields noticeable improvements over their original counterparts in Korea, India and Nigeria, indicating the effectiveness of retrieval adaptation.
+On cVQA, while Ravenea-CLIP generally exhibit large performance improvement compared to non-RAG baselines, certain countries, such as Indonesia, experience exhibit diminishing returns.
+Table 11: Ablation at cultural context. Models in gray are frozen retrievers, noisy characters, and unrelated documents. VLMs augmented with finetuned retriever generally perform better.
+Retriever Average DeepSeek-VL2-Tiny Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-78B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OV-7B $cVQA_{Accuracy\uparrow}$ W/O RAG 66.0 49.8 62.7 52.6 83.3 64.1 71.3 84.2 66.5 79.4 42.6 73.2 62.7 Unrelated documents 60.6 46.4 58.9 53.6 76.1 60.8 67.0 82.3 59.8 74.6 41.1 63.2 43.5 Random Characters 62.3 44.0 57.9 52.6 79.4 63.2 69.9 83.7 54.1 75.1 41.1 68.9 57.4 CLIP-L/14@224px 69.4 50.2 63.6 66.0 82.8 73.7 78.5 84.2 69.4 76.6 67.0 76.6 44.5 Ravenea-CLIP (ours) 71.2 50.7 64.1 69.9 82.3 76.1 81.3 85.2 69.9 79.9 69.4 75.1 50.2 $cIC_{RegionScore\uparrow}$ W/O RAG 54.4 26.5 53.1 69.4 71.4 38.8 49.0 65.3 73.5 75.5 28.6 53.1 49.0 Unrelated documents 48.8 18.4 46.9 57.1 61.2 28.6 51.0 63.3 59.2 71.4 24.5 51.0 53.1 Random Characters 51.2 24.5 44.9 61.2 67.3 36.7 44.9 67.3 61.2 73.5 28.6 59.2 44.9 CLIP-L/14@224px 64.1 38.8 59.2 71.4 69.4 65.3 65.3 71.4 73.5 73.5 61.2 69.4 51.0 Ravenea-CLIP (ours) 67.7 46.9 61.2 77.6 75.5 69.4 67.3 73.5 75.5 75.5 63.3 71.4 55.1
 Table 12: Four metrics comparison of RAG and Non-RAG methods for CCUB task.
-Open Weights Closed Weights Method Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OneVision-7B GPT-4.1 $CCUB_{Rouge-L}$ W/O RAG 18.3 18.7 18.2 21.8 18.2 16.5 17.8 19.6 19.8 16.7 15.9 20.6 15.4 21.4 19.6 Frozen Models SigLIP2-SO/14@384px 18.0 20.0 19.1 23.8 17.2 14.9 19.2 17.4 17.5 15.1 14.9 23.1 14.0 23.6 17.0 CLIP-L/14@224px 18.1 20.9 18.5 23.0 16.8 14.5 18.3 17.2 17.5 15.1 15.8 22.4 14.4 24.7 16.1 Finetuned Models VisualBERT 18.6 19.5 18.3 22.5 16.8 14.8 17.4 17.4 16.9 15.0 16.0 22.5 14.8 24.3 19.0 VL-T5 17.6 19.0 18.1 23.1 17.9 14.4 18.2 17.2 16.9 14.7 15.4 20.8 13.8 23.5 17.4 LLaVA-OneVision-7B 17.3 18.4 17.7 23.0 17.2 14.5 17.6 16.4 17.1 14.6 15.0 20.9 13.1 22.8 17.3 SigLIP2-SO/14@384px 18.4 19.3 19.2 24.5 18.5 14.8 18.4 18.2 17.2 15.2 16.4 22.2 14.3 23.5 18.9 CLIP-L/14@224px 18.1 19.6 18.6 23.9 17.1 14.7 18.0 17.3 17.7 16.0 16.3 21.5 13.7 24.2 17.0 $CCUB_{CIDER}$ W/O RAG 28.1 30.0 27.2 49.2 22.9 16.9 10.5 18.6 35.8 11.2 10.1 49.1 16.7 47.4 34.4 Frozen Models SigLIP2-SO/14@384px 25.2 39.2 20.5 58.5 16.9 7.6 12.4 7.1 6.3 4.7 6.1 62.0 10.1 51.4 17.2 CLIP-L/14@224px 23.8 40.9 18.8 59.6 13.0 6.2 9.6 7.8 4.8 4.7 6.6 56.3 11.1 49.7 14.6 Finetuned Models VisualBERT 26.1 40.2 24.0 60.9 16.3 8.2 10.0 7.5 5.1 6.4 10.2 51.1 15.1 49.5 24.4 VL-T5 24.5 37.0 20.6 51.1 16.0 7.3 10.2 5.0 4.0 6.8 7.5 52.5 10.3 48.4 21.6 LLaVA-OneVision-7B 21.7 38.8 18.5 62.6 13.2 4.8 8.4 5.7 3.4 5.7 4.5 48.9 11.4 42.2 17.6 SigLIP2-SO/14@384px 28.2 40.2 22.8 64.7 17.8 8.4 9.7 5.8 4.4 4.7 7.7 62.3 13.2 54.8 26.5 CLIP-L/14@224px 24.6 35.4 22.1 61.5 15.7 8.9 9.9 6.5 5.8 5.8 7.8 53.1 11.2 51.7 18.3 $CCUB_{BERTScore}$ W/O RAG 54.7 55.4 55.6 55.6 55.7 54.2 52.5 54.5 55.4 54.1 54.8 53.5 52.7 56.6 56.8 Frozen Models SigLIP2-SO/14@384px 54.9 57.2 56.4 57.9 55.2 53.0 55.1 54.8 55.0 52.4 54.0 56.9 52.6 58.7 54.8 CLIP-L/14@224px 54.8 57.6 56.2 57.7 54.8 52.4 54.8 55.1 55.3 52.6 53.7 55.9 52.5 59.4 54.6 Finetuned Models VisualBERT 54.8 57.1 56.5 57.6 54.9 52.8 54.6 54.9 54.4 52.1 53.8 55.7 52.8 58.5 56.2 VL-T5 54.2 56.7 56.0 58.1 54.9 52.6 53.9 54.0 54.4 52.0 53.4 54.4 52.1 57.8 55.7 LLaVA-OneVision-7B 54.2 56.6 55.5 58.2 54.5 52.1 53.9 54.2 54.6 51.6 53.3 55.0 51.8 58.3 54.9 SigLIP2-SO/14@384px 55.2 57.4 57.1 57.9 55.7 53.2 54.7 55.2 54.9 52.5 54.2 56.8 52.6 58.7 56.7 CLIP-L/14@224px 55.0 57.0 56.5 58.2 54.8 53.2 54.6 55.3 55.3 52.9 54.2 56.0 52.3 59.3 55.4 $CCUB_{CLIPScore}$ W/O RAG 19.1 19.1 18.8 19.6 18.3 18.2 18.4 18.7 19.2 18.4 18.4 20.2 18.5 20.2 18.7 Frozen Models SigLIP2-SO/14@384px 19.1 19.3 19.0 19.9 18.9 18.8 19.1 18.8 18.7 18.6 18.5 20.5 18.8 20.0 18.6 CLIP-L/14@224px 19.0 19.2 19.0 20.0 19.0 18.5 19.2 18.8 18.6 18.8 18.4 20.4 18.5 19.6 18.3 Finetuned Models VisualBERT 19.1 19.1 18.9 20.2 18.9 18.9 19.1 18.9 18.8 18.5 18.6 20.3 18.7 19.9 18.5 VL-T5 19.1 19.4 19.1 19.9 19.0 18.8 19.0 19.1 18.7 18.5 18.7 20.4 18.7 19.9 18.4 LLaVA-OneVision-7B 19.1 19.3 18.9 20.1 18.9 18.7 19.3 18.9 18.6 18.5 18.6 20.4 18.8 19.7 18.6 SigLIP2-SO/14@384px 19.1 19.2 18.9 20.1 18.8 18.6 19.1 19.0 18.5 18.6 18.5 20.4 18.8 19.8 18.5 CLIP-L/14@224px 19.0 19.2 18.9 20.0 19.1 18.6 19.2 18.9 18.7 18.6 18.4 20.1 18.8 19.7 18.6
-Figure 11: Average improvement across 14 VLMs in different countries with 4 retrievers for cVQA task.
-As shown in Figures 12 and 11, VLMs demonstrate varying degrees of performance shifts across cultural contexts in both the cIC and cVQA tasks.
-In the cIC setting, most fine-tuned retrievers yield noticeable improvements over their original counterparts, indicating the effectiveness of retrieval adaptation.
-Performance on the cVQA task reveals more nuanced outcomes.
-While fine-tuned retrievers generally exhibit large performance improvement compared to non-RAG baselines, certain countries, such as Spain and Indonesia, experience exhibit diminishing returns.
-These discrepancies may stem from the limited presence of culturally representative visual content in the training data.
-Although targeted image augmentation strategies were employed to alleviate this imbalance, the results suggest that data distribution remains a significant bottleneck.
-Understanding and addressing such cross-cultural performance disparities in multimodal tasks like cVQA will be an important direction for future work.
-Figure 12: Average improvement across 14 VLMs in different countries with 4 retrievers in cIC task.
-Appendix K Human Annotation & Evaluation Interface
-The annotation interface are shown in Figure 14 and 15. The interface for human evaluation in cIC task is shown in Figure 13.
-Figure 13: Human evaluation interface for cIC task.
-Figure 14: Human annotation instructions.
-(a) Annotation interface (1).
-(b) Annotation interface (2).
-Figure 15: Human annotation interfaces.
-Appendix L Correlation between Automatic Metrics and Human Judgments
-To assess the correlation between human preferences and automatic evaluation metrics, we compute Kendall’s Tau rank correlation.
-Human annotations are segmented according to the output chunks produced by corresponding 14 VLMs with each retriever variant (CaCLIP-based, CLIP-based, and non-RAG).
-Within each segment, we calculate the selection winning ratio for each retrieval method, yielding a human preference vector formed by concatenating these ratios across all evaluation instances.
-For the automatic evaluation, we extract BERTScore, CIDEr, ROUGE-L, and CLIPScore for each corresponding retrieval variant and VLM.
-These scores are similarly concatenated into a metric-based vector.
-Finally, we compute Kendall’s Tau between the human and metric vectors to quantify the consistency between automatic rankings and human judgments.
-Appendix M Prompts
+Method Average DeepSeek-VL2-Tiny DeepSeek-VL2 Qwen2.5-VL-3B Qwen2.5-VL-7B Qwen2.5-VL-72B InternVL3-2B InternVL3-8B InternVL3-38B Gemma3-4B Gemma3-27B Phi4-Multimodal Pixtral-12B LLaVA-OV-7B $CCUB_{Rouge-L}$ W/O RAG 18.0 19.4 18.1 24.3 18.3 17.1 17.4 18.6 17.0 15.3 14.4 15.2 15.1 23.6 Frozen Models SigLIP2-SO/14@384px 14.9 11.6 16.9 18.7 16.1 14.8 14.0 15.3 16.2 13.0 12.6 12.4 15.0 17.3 CLIP-L/14@224px 15.3 13.6 18.1 19.8 17.0 14.8 13.1 15.1 15.5 12.9 13.8 12.9 15.0 17.3 Finetuned Models VisualBERT 14.6 11.1 16.1 17.4 15.7 14.7 12.2 15.3 15.6 12.9 13.7 11.8 15.8 18.0 VL-T5 14.4 10.6 15.9 18.0 16.8 14.3 13.0 15.1 15.6 12.6 13.1 10.1 14.8 17.2 LLaVA-OV-7B 14.0 10.0 15.5 17.9 16.1 14.4 12.4 14.3 15.8 12.5 12.7 10.2 14.1 16.5 Ravenea-SigLIP 14.8 11.7 16.4 17.3 15.5 15.5 13.3 16.1 15.6 13.2 13.0 13.0 15.1 16.9 Ravenea-CLIP 15.3 12.1 16.7 19.5 18.1 15.6 12.8 15.6 16.1 12.5 13.3 13.6 14.7 18.3 $CCUB_{CIDER}$ W/O RAG 22.2 37.4 26.1 60.7 19.1 11.0 5.7 20.2 10.8 7.5 6.5 6.3 18.0 59.2 Frozen Models SigLIP2-SO/14@384px 8.5 18.3 12.5 17.5 9.8 3.5 2.0 5.6 3.0 1.4 2.0 0.6 13.1 20.5 CLIP-L/14@224px 9.0 21.5 17.0 18.9 11.5 4.3 0.6 3.2 3.3 3.4 2.7 1.6 10.6 18.0 Finetuned Models VisualBERT 9.5 19.3 16.0 19.9 9.2 4.1 0.5 6.0 1.8 3.1 6.1 1.3 18.1 18.6 VL-T5 7.2 16.1 12.6 10.1 8.9 3.2 0.2 3.5 0.7 3.5 3.4 0.9 13.3 17.5 LLaVA-OV-7B 7.1 17.9 10.5 21.6 6.1 0.7 1.2 4.2 0.1 2.4 0.4 1.2 14.4 11.3 Ravenea-SigLIP 8.0 16.8 12.9 12.5 10.0 5.4 0.4 5.3 3.6 1.8 1.8 0.7 13.3 19.0 Ravenea-CLIP 8.5 16.8 11.8 17.3 10.7 5.0 0.1 4.5 3.5 1.3 1.9 1.1 12.6 23.3 $CCUB_{BERTScore}$ W/O RAG 53.6 53.3 55.3 56.2 54.5 53.2 51.7 54.2 53.6 50.1 51.4 51.9 53.5 57.5 Frozen Models SigLIP2-SO/14@384px 50.9 46.7 52.3 54.2 53.1 50.4 48.9 51.8 51.9 48.1 50.3 49.3 51.2 53.7 CLIP-L/14@224px 51.9 49.7 53.1 54.5 53.7 51.6 49.4 52.8 52.5 49.2 50.9 49.8 52.7 55.0 Finetuned Models VisualBERT 50.6 46.6 52.4 53.9 52.8 50.2 48.4 51.9 51.3 47.8 50.1 48.1 51.4 53.5 VL-T5 50.2 46.2 51.9 54.4 52.8 50.0 47.7 51.0 51.3 47.7 49.7 46.8 50.7 52.8 LLaVA-OV-7B 50.2 46.1 51.4 54.5 52.4 49.5 47.7 51.2 51.5 47.3 49.6 47.4 50.4 53.3 Ravenea-SigLIP 51.2 46.7 53.0 52.7 52.8 51.7 49.1 53.0 52.6 48.9 50.5 49.3 51.6 54.1 Ravenea-CLIP 52.5 49.2 54.2 54.9 54.2 52.9 49.3 53.2 53.0 49.0 51.4 51.3 53.4 56.4 $CCUB_{CLIPScore}$ W/O RAG 33.4 33.3 34.1 32.5 33.8 33.4 34.1 34.3 34.5 33.9 33.8 32.4 32.6 32.2 Frozen Models SigLIP2-SO/14@384px 32.7 29.6 32.9 33.4 32.7 33.0 33.8 33.0 33.7 32.1 33.7 31.8 31.9 33.3 CLIP-L/14@224px 32.7 31.4 33.4 33.7 33.2 33.7 35.1 34.1 34.1 33.2 34.1 32.7 32.5 33.6 Finetuned Models VisualBERT 32.7 29.4 32.8 33.7 32.7 33.1 33.8 33.1 33.8 32.0 33.8 31.6 31.8 33.2 VL-T5 32.8 29.7 33.0 33.4 32.8 33.0 33.7 33.3 33.7 32.0 33.9 31.7 31.8 33.2 LLaVA-OV-7B 32.7 29.6 32.8 33.6 32.7 32.9 34.0 33.1 33.6 32.0 33.8 31.7 31.9 33.0 Ravenea-SigLIP 32.7 29.8 33.0 32.1 33.1 33.5 33.8 33.1 33.8 32.7 33.7 31.7 32.0 33.1 Ravenea-CLIP 33.5 31.6 33.4 34.0 33.6 33.7 34.7 34.2 34.2 33.0 33.9 32.8 32.8 33.3
+Appendix K Prompts
 The prompts that we used to collect the documents and two downstream tasks were designed to ensure consistency across different models.
 One prompt example for cultural caption generation is shown in Table 13.
-The prompts for both with and without retrieval augmentation for VLMs are shown in Table 12 and 11.
+The prompts for both with and without retrieval augmentation for VLMs are shown in Table 15 and 14.
 Table 13: Prompt example for generating captions used to retrieve the Wiki documents.
-Prompt for data collection GPT-4o SYSTEM: Default
-USER:
-Generate a culture related caption given the image with around 2 sentences.
+Prompt for data collection GPT-4o SYSTEM: Default USER:[Uncaptioned image] Generate a culture related caption given the image with around 2 sentences.
 Please include the name of the thing shown in the picture within the caption if it’s strongly related to the culture.
@@
 USER:
-Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
+Answer the following multiple choice question for the image. The last line must be of the following format: ’Answer: \$LETTER’ (without quotes) where LETTER must be one of A, B, C, or D.
 Question: Which city is famous for such artwork?
@@
 Write a concise, one-sentence caption for the given image.
-The generated caption must contain the visual content and culturally relevant elements of the image.
-Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph...").
-Do not generate multiple options.
+The generated caption must contain the visual content, culturally-relevant elements, and the country’s name.
+Avoid phrases like ’This image’ or ’The image’.
 Table 15: Prompt examples with RAG. Multimodal prompt samples with interleaved image are shown for both CVQA and CCUB tasks.
@@
 USER:
-Answer the following multiple choice question. The last line of your response must be of the following format: Ánswer: \$LETTER(́without quotes) where LETTER is one of ABCD.
-The scope of the question is strictly limited to the given image. However, please analyze and incorporate information from both the image and the following document to answer the question.
+Answer the following multiple choice question for the image. The last line must be of the following format: ’Answer: \$LETTER’ (without quotes) where LETTER must be one of A, B, C, or D.
+Use the shared culture between the image and the following document to answer the question.
 Document:
@@
 Write a concise, one-sentence caption for the given image.
-The generated caption must contain the visual content and culturally relevant elements of the image.
-Avoid explicit references to the image itself (e.g., "This image shows...", "Pictured here is...", "In this photograph...").
-Do not generate multiple options.
-Please consider the following context:
+The generated caption must contain the visual content, culturally-relevant elements, and the country’s name.
+Avoid phrases like ’This image’ or ’The image’.
+Please consider the following DOCUMENT as supplementary material for the image.
+Mention the country’s name of the culture.
+Document:
 Architecture of Nigeria was historically influenced by environmental conditions as well as social and cultural factors.
@@
 A one or two storey timber house building made with pre-fabricated material components and designed with the influence of classic antiquity styles served as mission house for the missionaries...
+Appendix L Human Annotation & Evaluation Interface
+The annotation interfaces are shown in Figure 13 and 14. The interface for human evaluation in cIC task is shown in Figure 12.
+Figure 12: Human evaluation interface for cIC task.
+Figure 13: Human annotation instructions.
+(a) Annotation interface (1).
+(b) Annotation interface (2).
+Figure 14: Human annotation interfaces.